### v1

# Guido Franchetti1 and Derek Harland2

Abstract
It is well-known that the $L^{2}$ metric on the moduli space of hyperbolic monopoles, defined using the Coulomb gauge-fixing condition, diverges. This article shows that an alternative gauge-fixing condition inspired by supersymmetry cures this divergence. The resulting geometry is a hyperbolic analogue of the hyperkähler geometry of Euclidean monopole moduli spaces.
1 Introduction
Moduli spaces and their metrics play a central role in the study of monopoles. The $L^{2}$ metric on the moduli space of monopoles in Euclidean space was introduced by Manton [23], who proposed that the geodesics of this metric should approximate the dynamics of slowly-moving monopoles. This conjecture was later confirmed by Stuart [28]. Inspired by Manton’s proposal, Atiyah–Hitchin calculated the metric on the moduli space of centred 2-monopoles [2]. Their calculation exploited the fact that the moduli space metrics are hyperkähler, and the Atiyah–Hitchin metric is now one of the best-known examples of a hyperkähler manifold. Many other prominent examples of hyperkähler metrics arise as moduli spaces of monopoles [22, 8, 9, 10]. The moduli space metric can be used to study not only classical but also quantum mechanics of monopoles [15]. Famously, Sen [25] made predictions for the $L^{2}$ cohomology of centred monopole moduli spaces, and these conjectures have been confirmed for charges 2 [18] and 3 [20].
Monopoles on hyperbolic space share much mathematical structure with their Euclidean counterparts, including correspondences with spectral curves, rational maps, and (discrete) Nahm equations, so it would be natural to expect their moduli spaces to admit interesting metrics. However, it has been known for a long time that the natural $L^{2}$ metric on the hyperbolic monopole moduli space diverges [6]. In spite of this difficulty, many attempts have been made to discover natural geometric structures on their moduli spaces. Hitchin [17] constructed an Einstein metric on the moduli space of inversion-symmetric charge 2 monopoles using twistor theory. Unfortunately, it has proven difficult to generalise this approach to higher charges [19], and it is not clear what relevance this metric has to monopole dynamics. More recently, Bielawski–Schwachhofer [5, 4] introduced pluricomplex structures on the moduli spaces, building on twistor theory developed by Nash [24]. These structures also arose from an analysis of the supersymmetry of hyperbolic monopoles [11]. As with Hitchin’s work, the relevance of Bielawski–Schwachhofer’s approach to classical dynamics of monopoles is unclear because it is formulated in terms twistor theory.
Gibbons–Warnick [16] and Franchetti–Ross [12] analysed the dynamics of well-separated monopoles, starting from a lagrangian for charged point particles. Interestingly, the metrics that they obtained were the LeBrun metrics, which had been studied much earlier in an entirely different context [21]. Franchetti–Schroers [13] calculated the $L^{2}$ metric for a circle-invariant 1-instanton on $S^{4}$, giving a finite metric on the moduli space of hyperbolic 1-monopoles which however is invariant only under a rotational subgroup of the $H^{3}$ isometry group. Finally, Sutcliffe [29] proposed a way to renormalise the divergent integral that defines the $L^{2}$ metric on the hyperbolic monopole moduli space. His boundary metric is related to the $L^{2}$ metric on the space of abelian connections on $S^{2}$, which had been considered earlier by Braam-Austin [7]. This metric was calculated explicitly in the case of inversion-symmetric charge 2 monopoles [30]. The asymptotic geometry of the boundary metric is very different from the metric of Franchetti–Ross [12, 14], so it is difficult to reconcile Sutcliffe’s metric with other approaches to monopole dynamics. It is also not clear that the boundary metric admits any analogue of the hyperkähler property.
This article introduces a new approach to the geometry of hyperbolic monopoles. The starting point is a gauge-fixing condition introduced by Figueroa-O’Farrill–Gharamti [11]. On hyperbolic space with sectional curvature $-s^{2}$ it takes the form
$\ast d^{A}\ast a+[\Phi,\phi]\pm{2i}{s}\phi=0.$ (1)
In this equation $a$ and $\phi$ are a Lie-algebra valued 1-form and 0-form describing a tangent vector to the moduli space. In the limit $s=0$ it reduces to the standard Coulomb gauge-fixing condition used to define the hyperkähler metric for Euclidean monopoles. This gauge-fixing condition was motivated by supersymmetry: it allowed the bosonic and fermionic zero-modes of the monopole to be related by supersymmetry transformations.
In this article we show that the gauge-fixing condition (1), subject to suitable boundary conditions, admits a unique solution. In other words, every tangent vector to the moduli space has a unique representative solving (1). We moreover show that the $L^{2}$ inner product of any two such tangent vectors is finite and so gives pointwise bilinear forms $g_{\pm}$ on the tangent bundle of the moduli space of hyperbolic monopoles. Thus the gauge fixing condition (1) solves the problem of the divergent $L^{2}$ metric.
The gauge-fixing condition (1) is unusual in that it explicitly involves the complex number $i$. So the solution $(a,\phi)$ should be expected to take values in a complex, rather than a real, Lie algebra. It is this fact that allows (1) to evade the divergence that afflicts more conventional approaches to $L^{2}$ metrics. However, this strength is also potentially a weakness: if $a$ and $\phi$ are not real then there is no reason to expect the integral of their square to be real. We nevertheless show that $g_{\pm}$ are real in at least two cases: the moduli space of charge 1 monopoles and the moduli space of inversion-symmetric charge $n$ monopoles. In particular, the hyperbolic analogue of the Atiyah–Hitchin metric is real.
We will moreover show that the gauge-fixing condition (1) leads to a natural decomposition
$T^{\mathbb{C}}M\cong E^{\pm}\otimes K^{\pm}$ (2)
of the complexified tangent bundle as a complex vector bundle $E^{\pm}$ and a 2-dimensional complex vector space $K^{\pm}$. This decomposition encodes the pluricomplex structures of Bielawski–Schwachhofer [5] and allows us to show that they are compatible with bilinear forms $g_{\pm}$. So $g_{\pm}$ appear to be natural hyperbolic counterparts to the hyperkähler metrics on moduli spaces of Euclidean monopoles. They also seem to be consistent with a physical interpretation of monopoles as particles whose inertial mass is equal to their energy. This is confirmed by an explicit calculation of metric components associated with Killing vector fields.
Throughout this article we assume that the mass parameter $p\in\mathbb{R}_{>0}$ of the monopole is a half-integer. This has two benefits. First, it provides natural boundary conditions for (1) by identifying hyperbolic monopoles with circle-invariant instantons on $S^{4}$ [1]. And second, it allows for streamlined analytic proofs that exploit the compactness of $S^{4}$. Relaxing the assumption $2p\in\mathbb{N}$ would result in lengthier proofs involving more sophisticated analysis, which we leave for future work.
This article is organised as follows. Section 2 defines the hyperbolic monopole moduli spaces and states our main results for existence and uniqueness of solutions to (1) and finiteness of the $L^{2}$ integrals $g_{\pm}$, Theorems 3 and 4. Section 3 is devoted to the proof of Theorem 3. Section 4 establishes the decomposition (2) and explains the compatibility between $g_{\pm}$ and the pluricomplex structures. It also completes the proof of Theorem 4. Finally, section 5 contains some explicit calculations of $g_{\pm}$ in particular settings.
2 The monopole metric
2.1 Preliminaries
A hyperbolic monopole is a pair $(A,\Phi)$ consisting of a connection $A$ and endomorphism $\Phi$ on a trivial vector bundle $V$ over hyperbolic space $H^{3}$ with sectional curvature $-1$. These are required to satisfy the Bogomolny equation,
$d^{A}\Phi=\ast F^{A},$ (3)
together with some boundary conditions to be specified. In this paper attention will be restricted to gauge group $SU(2)$, so that $V$ has rank 2, $\Phi$ is an $\mathfrak{su}(2)$-valued function, and $A$ is an $\mathfrak{su}(2)$-valued 1-form. In this situation the boundary conditions require that:
  • $V$, $\Phi$ and $A$ extend to the boundary $S^{2}_{\infty}$ of $H^{3}$;
  • $V|_{S^{2}_{\infty}}$ decomposes as a sum of line bundles $L^{\pm}$ with first Chern number $c_{1}(L^{\pm})=\pm n$ for some $n\in\mathbb{N}$;
  • the bundles $L^{\pm}$ are eigenbundles of $\Phi_{S^{2}_{\infty}}$ with eigenvalues $\mp ip$ for some $p>0$;
  • the connection $A|_{S^{2}_{\infty}}$ is a sum of $U(1)$ connections on these line bundles.
The $SU(2)$-structure means that $L^{-}$ is naturally isomorphic to the dual of $L^{+}$, and that the connection on $L^{-}$ is the dual of the connection on $L^{+}$. The integer $n$ is known as the charge of the monopole, while $p$ is known as the mass parameter.
Solutions of the Bogomolny equation (3) minimise the energy,
$E=-\frac{1}{4}\int_{H^{3}}\operatorname{Tr}(d^{A}\Phi\wedge\ast d^{A}\Phi+F^{A}\wedge\ast F^{A}).$ (4)
This fact follows from a simple integration by parts argument:
$E$ $=-\frac{1}{4}\int_{H^{3}}\operatorname{Tr}(d^{A}\Phi-\ast F^{A})\wedge\ast(d^{A}\Phi-\ast F^{A})-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(d^{A}\Phi\wedge F^{A})$ (5)
$\geq 0-\frac{1}{2}\int_{S^{2}_{\infty}}\operatorname{Tr}(\Phi F^{A})=2\pi np.$ (6)
[…]
In the particular case $2p\in\mathbb{N}$, charge $n$ hyperbolic monopoles coincide with circle-invariant instantons on $S^{4}$ with second Chern number $2np$ [1]. This follows from the fact that the metric on $S^{1}\times H^{3}$ is conformally equivalent to the metric on $S^{4}\setminus S^{2}$, with the circle action on $S^{1}\times H^{3}$ corresponding to rotations of $S^{4}$ that fix the removed $S^{2}$. Hyperbolic monopoles can be identified with circle-invariant anti-self-dual connections $A+\Phi\,d\theta$ on $S^{1}\times H^{3}$. Since the anti-self-dual equations are conformally invariant, any circle-invariant instanton on $S^{4}$ gives rise to a solution of the Bogomolny equation on $S^{1}\times H^{3}$.
The conformal equivalence of $S^{1}\times H^{3}$ and $S^{4}\setminus S^{2}$ can be realised explicitly as follows. We model $S^{4}$ as the set of unit vectors in $\mathbb{R}^{5}$. By writing $y_{1}=r_{1}\cos\theta$, $y_{2}=r_{1}\sin\theta$ with $r_{1}>0$, the metric on $\mathbb{R}^{5}\setminus\mathbb{R}^{3}$ is seen to be conformally equivalent to the metric on $S^{1}\times H^{4}$:
$\frac{1}{r_{1}^{2}}\sum_{\mu=1}^{5}dy_{\mu}dy_{\mu}=d\theta^{2}+\frac{1}{r_{1}^{2}}(dr_{1}^{2}+dy_{3}^{2}+dy_{4}^{2}+dy_{5}^{2}).$ (7)
The equation defining $S^{4}$, $r_{1}^{2}+y_{3}^{2}+y_{4}^{2}+y_{5}^{2}=1$, determines an embedding $H^{3}\subset H^{4}$. We thus obtain a conformal identification of $S^{4}\setminus S^{2}$ with $S^{1}\times H^{3}$. The circle group acts naturally on both. The excluded set $S^{2}\subset S^{4}$, corresponding to $r_{1}=0$, is the fixed point set of the circle action and is naturally identified with the boundary of $H^{3}$.
A circle-invariant instanton $B$ on $S^{4}$ pulls back to a circle-invariant anti-self-dual connection on $S^{1}\times H^{3}$. One can find a gauge transformation $g:S^{1}\times H^{3}\to SU(2)$ such that
$g^{-1}Bg+g^{-1}dg=\Phi\,d\theta+A,$ (8)
with $(A,\Phi)$ a monopole on $H^{3}$, and $\theta$ a coordinate on $S^{1}$. Since $B$, $\Phi$ and $A$ are all invariant under $\theta\mapsto\theta+2\pi$, $g(\theta+2\pi)g(\theta)^{-1}$ is parallel with respect to $B$. Since $B$ is irreducible, this means that $g(\theta+2\pi)g(\theta)^{-1}=\pm I_{2}$. The holonomy of $B$ about $S^{1}$ is therefore given by ±$\pm\exp(-2\pi\Phi)$. Since the holonomy tends to $\mathrm{Id}$ as one approaches the fixed $S^{2}\subset S^{4}$, the eigenvalues of $\Phi$ approach ±$\pm ip$ for some $p\in\tfrac{1}{2}\mathbb{N}$. Thus circle-invariant instantons give rise to hyperbolic monopoles with half-integer mass parameter $p$; the converse to this statement was proved in [26, 27].
For later use, we now describe the circle action on the bundle $V\to S^{4}$ in more detail. Over $S^{1}\times H^{3}$, we have a trivialisation $V=S^{1}\times H^{3}\times\mathbb{C}^{2}$ and the circle action is $(\theta,x,s)\mapsto(\theta+t,x,s)$. The gauge transformation $g$ in (8) must take the form
$g=\begin{pmatrix}e^{-i\theta p}&0\\ 0&e^{i\theta p}\end{pmatrix}$ (9)
with respect to the splitting $V=L^{+}\oplus L^{-}$ over $S^{2}_{\infty}$, because $B(\frac{\partial}{\partial\theta})\to 0$ as one approaches this fixed $S^{2}$. Therefore, after making a gauge transformation with $g$, the circle action on the bundle $L^{+}\oplus L^{-}$ over $S^{1}\times S^{2}_{\infty}$ is
$(\theta,x,s_{+},s_{-})\mapsto(\theta+t,x,e^{-itp}s_{+},e^{itp}s_{-}).$ (10)
Therefore the circle group acts with weights pminus-or-$\mp p$ on the bundles $L^{\pm}$ over the $S^{2}\subset S^{4}$ that is fixed by the circle action. Strictly speaking, if $2p$ is odd the group that acts on $V\to S^{4}$ is a double cover of the circle group.
2.2 The framed and unframed moduli spaces
The Bogomolny equation is invariant under gauge transformations, and it is natural to study the moduli space of solutions modulo gauge transformations. In fact, there are two natural moduli spaces to consider, the framed and unframed moduli spaces. We define an (unframed) gauge transformation to be a function $g:H^{3}\to SU(2)$ that extends to a smooth circle-invariant section of the bundle $\operatorname{Aut}(V)\to S^{4}$. Note that circle-invariance demands that on the fixed $S^{2}\subset S^{4}$, $g$ maps $L^{+}$ to $L^{+}$ and $L^{-}$ to $L^{-}$. The moduli space of unframed monopoles is the space $M^{u}_{n,p}$ of monopoles with charge $n$ and mass parameter $p$ that extend smoothly to $S^{4}$, modulo unframed gauge transformations. This is known to be a smooth manifold of dimension $14n-$ 1 [7].
To define framed monopoles, we need to recall the definition of the Chern–Simons number. Let $A^{0},A^{1}$ be two connections on $H^{3}$ and let $\tilde{A}=(1-t)A^{0}+tA^{1}$ be a connection on $[0,1]\times H^{3}$. The Chern-Simons number of $A_{0},A_{1}$ is defined to be
$CS[A^{0},A^{1}]=2\int_{[0,1]\times H^{3}}\operatorname{Tr}(F^{\tilde{A}}\wedge F^{\tilde{A}}).$ (11)
By explicitly evaluating the integral in $t$, one can obtain a more explicit formula for the Chern-Simons number:
$CS[A^{0},A^{1}]=\int_{H^{3}}\operatorname{Tr}\left(A^{1}\wedge dA^{1}+\frac{2}{3}A^{1}\wedge A^{1}\wedge A^{1}\right)\\ -\int_{H^{3}}\operatorname{Tr}\left(A^{0}\wedge dA^{0}+\frac{2}{3}A^{0}\wedge A^{0}\wedge A^{0}\right)+\int_{S^{2}_{\infty}}\operatorname{Tr}(A^{0}\wedge A^{1}).$ (12)
Lemma 1.
Let $A^{0},A^{1},A^{2}$ be connections satisfying the boundary conditions of a monopole, and let $g:H^{3}\to SU(2)$ be a gauge transformation. Then $CS[A^{0},A^{1}]=CS[g^{-1}A^{0}g+g^{-1}dg,g^{-1}A^{1}g+g^{-1}dg]$. Moreover,
$CS[A^{0},A^{2}]=CS[A^{0},A^{1}]+CS[A^{1},A^{2}],\quad CS[A^{1},A^{0}]=-CS[A^{0},A^{1}]$ (13)
Proof.
Gauge invariance is immediate from the definition (11), as $F^{\tilde{A}}\mapsto g^{-1}F^{\tilde{A}}g$ under gauge transformations. The first condition of (13) can be proved from the explicit formula (12):
$\begin{split}CS[A^{0},A^{1}]+&CS[A^{1},A^{2}]-CS[A^{0},A^{2}]\\ &=\int_{S^{2}_{\infty}}\operatorname{Tr}(A^{0}\wedge A^{1}+A^{1}\wedge A^{2}-A^{0}\wedge A^{2})\\ &=\int_{S^{2}_{\infty}}\operatorname{Tr}(A^{1}\wedge A^{1}+(A^{0}-A^{1})\wedge(A^{1}-A^{2})).\end{split}$ (14)
The first of these two terms vanishes by cyclicity of the trace and skew-symmetry of the wedge product. The second vanishes because $A^{i}-A^{j}$ respects the splitting $V_{S^{2}_{\infty}}=L^{+}\oplus L^{-}$, in other words, it take values in an abelian subalgebra $\mathfrak{u}(1)\subset\mathfrak{su}(2)$. Thus the first identity of (13) holds. The second is immediate from either equation (11) or (12). ∎
We say that two connections $A^{1},A^{2}$ are framed gauge equivalent if they are gauge equivalent and $CS[A^{1},A^{2}]=0$. The moduli space of framed monopoles is the space $M^{f}_{n,p}$ of monopoles with charge $n$ and mass parameter $p$ that extend smoothly to $S^{4}$, modulo framed gauge equivalence.
There is a natural map $M^{f}_{n,p}\to M^{u}_{n,p}$ which sends a framed monopole to its (unframed) gauge equivalence class. Give a choice of reference monopole $(A^{0},\Phi^{0})$, there is also a natural map $M^{f}_{n,p}\to\mathbb{R}$ which sends $(A,\Phi)$ to $CS[A^{0},A]$. Together these give a map
$M^{f}_{n,p}\to\mathbb{R}\times M^{u}_{n,p}$ (15)
which we claim is a bijection. The map is injective because, by definition, if $(A^{1},\Phi^{1})$ and $(A^{2},\Phi^{2})$ are gauge equivalent and $CS[A^{0},A^{1}]=CS[A^{0},A^{2}]$ they are framed gauge equivalent by equation (13). To show that the map is surjective we use the following lemma, which also shows that the unframed gauge transformations are the ones generated by the Higgs field:
Lemma 2.
Let $(A,\Phi)$ be a monopole, let $g_{s}=\exp(s\Phi)$ and let $A^{s}=g_{s}^{-1}dg_{s}+g_{s}^{-1}Ag_{s}$. Then
$CS[A,A^{s}]=-4\pi nps.$ (16)
Proof.
We prove the identity by direct calculation from the explicit formula (12) (one can also prove this directly from the definition (11)). We find that $\frac{d}{ds}A_{s}=d^{A_{s}}\Phi$ and hence that
$\begin{split}\frac{d}{ds}CS[A,A_{s}]&=\int_{H^{3}}\operatorname{Tr}(d^{A_{s}}\Phi\wedge dA_{s}+A_{s}\wedge d(d^{A_{s}}\Phi)+2A_{s}\wedge A_{s}\wedge d^{A_{s}}\Phi)\\ &\quad+\int_{S^{2}_{\infty}}\operatorname{Tr}(A\wedge d^{A_{s}}\Phi)\\ &=\int_{H^{3}}\big{[}2\operatorname{Tr}(d^{A_{s}}\Phi\wedge F^{A_{s}})-d\operatorname{Tr}(A_{s}\wedge d^{A_{s}}\Phi)\big{]}\\ &\quad+\int_{S^{2}_{\infty}}\operatorname{Tr}(A\wedge d^{A_{s}}\Phi)\\ &=-4\pi np.\end{split}$ (17)
In going from the second to the final line we applied Stokes’ theorem, used the fact that $d^{A^{s}}\Phi|_{S^{2}_{\infty}}=0$ since the monopole abelianises at the boundary, and evaluated the remaining term using the identity (6). The result follows. ∎
To prove surjectivity of (15), let $t\in\mathbb{R}$ and let $(A,\Phi)$ be a monopole, thought of as an element of $M^{u}_{n,p}$. We need to find a gauge equivalent monopole whose Chern-Simons number is $t$. Let
$s=\frac{1}{4\pi np}(CS[A^{0},A]-t)$ (18)
Lemma 2, combined with equation (13), shows that gauge-transforming $(A,\Phi)$ with $\exp(s\Phi)$ produces a monopole $\tilde{A}$ with $CS[A^{0},\tilde{A}]=t$.
Our definition of the framed moduli space is different from others in the literature, and we explain the relation with other definitions in Appendix A. Our definition is motivated by monopole dynamics. One can extend the energy functional (4) to define a lagrangian gauge theory on ×$\mathbb{R}\times H^{3}$. Dynamical configurations consist of a gauge field $A$ and scalar field $\Phi$ on ×$\mathbb{R}\times H^{3}$. Then, for fixed $t_{0},t_{1}\in\mathbb{R}$, the quantity
$\Theta(t_{0},t_{1})=2\int_{[t_{0},t_{1}]\times H^{3}}\operatorname{Tr}(F^{A}\wedge F^{A})$ (19)
is clearly gauge-invariant and equal to 0 for any static configuration. By working in a gauge where $A_{t}=0$, slowly moving dynamical configurations can be approximated by paths in the monopole moduli space, parametrised by $t\in\mathbb{R}$. Consider a path of the form $(A^{t},\Phi^{t})=(g_{t}^{-1}Ag_{t}+g_{t}^{-1}dg_{t},g_{t}^{-1}\Phi g_{t})$, where $(A,\Phi)$ is a fixed monopole on $H^{3}$ and $g$ is a $t$-dependent gauge transformation. It can be shown that $\Theta(s,t)=CS[A^{s},A^{t}]$. In particular, Lemma 2 shows that, for the choice $g_{t}=\exp(t\Phi)$, $\Theta$ is non-zero, even though this path is trivial in the unframed moduli space. This example shows that in order to describe monopole dynamics it is necessary to work in the framed moduli space, rather than the unframed moduli space.
Paths in the moduli space generated by gauge transformation with $\Phi$ are known as dyons and carry an electric charge, which is defined to be the abelian electric flux through $S^{2}_{\infty}$ and computed by the integral $\frac{1}{2p}\int_{S^{2}_{\infty}}\operatorname{Tr}(\Phi\ast\frac{dA}{dt})$. If $\frac{dA}{dt}=cd^{A}\Phi$ then the electric flux is $\frac{c}{2p}\int_{S^{2}_{\infty}}\operatorname{Tr}(\Phi F^{A})$ and the rate of change of the Chern-Simons number is $4c\int_{H^{3}}\operatorname{Tr}(d^{A}\Phi\wedge F^{A})$. Integration by parts then shows that the electric charge is proportional to the rate of change of the Chern-Simons number.
2.3 The moduli space metric
A tangent vector to the framed moduli space at $(A,\Phi)$ is described by an $\mathfrak{su}(2)$-valued 1-form $a$ and 0-form $\phi$ solving the linearised Bogomolny equations,
$d^{A}\phi+[a,\Phi]=\ast d^{A}a,$ (20)
such that $a+\phi\,d\theta$ extends to a smooth circle-invariant section of $\Lambda^{1}(S^{4})\otimes\operatorname{End}(V)$. Given any $\mathfrak{su}(2)$-valued function $\chi$ that extends to a circle-invariant section over $S^{4}$, $(a,\phi)=(d^{A}\chi,[\Phi,\chi])$ is a solution of (20) representing an infinitesimal gauge transformation. A short calculation shows that if $g_{s}$ is a family of gauge transformations such that $g_{0}=0$, $g_{s}^{-1}\partial_{s}g_{s}|_{s=0}=\chi$ then
$\left.\frac{d}{ds}\right|_{s=0}CS[A,g_{s}^{-1}Ag_{s}+g_{s}^{-1}dg_{s}]=2\int_{H^{3}}\operatorname{Tr}(F^{A}\wedge d^{A}\chi).$ (21)
Thus we define the space $G^{u}_{A,\Phi}$ of infinitessimal unframed gauge transformations to be the set of all pairs $(d^{A}\chi,[\Phi,\chi])$ such that $\chi:H^{3}\to\mathfrak{su}(2)$ extends smoothly to a circle-invariant section of $\operatorname{End}(V)\to S^{4}$. We define the space $G^{f}_{A,\Phi}$ of infinitessimal framed gauge transformations to be the subspace such that $\int_{H^{3}}\operatorname{Tr}(F^{A}\wedge d^{A}\chi)=0$. These are both subspaces of the space $T_{A,\Phi}C_{n,p}$ of pairs $(a,\phi)$ that extend smoothly to $S^{4}$ in a circle-invariant way and satisfy equation (20). The tangent space to the framed and unframed moduli spaces are the quotient spaces
$T_{A,\Phi}M^{f}_{n,p}=T_{A,\Phi}C_{n,p}/G^{f}_{A,\Phi},\quad T_{A,\Phi}M^{u}_{n,p}=T_{A,\Phi}C_{n,p}/G^{u}_{A,\Phi}.$ (22)
An $L^{2}$ metric on the moduli space is a metric obtained by integrating $X=(a,\phi)$:
$g(X,X)=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(a\wedge\ast a+\phi\ast\phi).$ (23)
In order to make this well-defined, we must choose a particular representative $(a,\phi)$ of a gauge equivalence class; in other words, we must fix a gauge. Unfortunately, any real representative of a non-zero tangent vector results in a divergent integral. This is because unframed hyperbolic monopoles have the holographic property of being determined by the abelian connection induced at infinity [7]. Any representative $a$ of a non-zero tangent vector must be non-zero when restricted to the line bundles $L^{\pm}$ over $S^{2}_{\infty}$. It follows that $a^{\prime}=-\operatorname{Tr}(a\Phi)/\sqrt{-\operatorname{Tr}(\Phi^{2})}$ gives a non-vanishing 1-form on any geodesic 2-sphere of sufficiently large radius. The metric $dr^{2}+\sinh^{2}(r)g_{S^{2}}$ of $H^{3}$ is such that the 1-form $a^{\prime}$ on $S^{2}$ satisfies $\ast a^{\prime}=\ast_{S^{2}}a^{\prime}\wedge dr$. Thus there exist positive constants $R,C$ such that
$g(X,X)\geq\frac{1}{2}\int_{r\geq R}a^{\prime}\wedge\ast_{S^{2}}a^{\prime}\wedge dr\geq\int_{R}^{\infty}C\,dr>\infty.$ (24)
Thus the integral determining the metric is divergent.
One way to avoid this divergence is to choose a non-real representative of a real tangent vector. Using an $\mathfrak{sl}(2,\mathbb{C})$-valued gauge transformation, we can choose a representative $(a,\phi)$ that is not antihermitian, but is such that $(a+a^{\dagger},\phi+\phi^{\dagger})$ is pure gauge. The advantage of doing so is that it opens up the possibility of $a|_{S^{2}_{\infty}}$ being a holomorphic (or antiholomorphic) 1-form. A 1-form $a_{\infty}$ that is of type $(1,0)$ or $(0,1)$ with respect to the complex structure of $S^{2}_{\infty}$ satisfies $a_{\infty}\wedge\ast_{S^{2}}a_{\infty}=0$. Thus the divergence of the $L^{2}$ metric could be avoided if the restriction of $a$ to $S^{2}_{\infty}$ were holomorphic or antiholomorphic.
A non-real gauge-fixing condition has been proposed by Gharamti–Figueroa-O’Farrill [11]. It takes the form
$\ast d^{A}\ast a_{\pm}+[\Phi,\phi_{\pm}]\pm 2i\phi_{\pm}=0.$ (25)
It differs from the more familiar harmonic gauge-fixing condition used in [7, 30] by the term ±$\pm 2i\phi_{\pm}$. Equation (25) is not one but two gauge-fixing conditions corresponding to the choice of sign. The two representatives $(a_{\pm},\phi_{\pm})$ of a real tangent vector to the moduli space will differ by a gauge transformation. We are free to choose $(a_{+},\phi_{+})=(-a_{-}^{\dagger},-\phi_{-}^{\dagger})$, as this choice clearly solves the “+” gauge fixing condition if $(a_{-},\phi_{-})$ solves the “--” condition, and, if $(a_{-},\phi_{-})$ represents a real tangent vector, it differs from $(-a_{+}^{\dagger},-\phi_{+}^{\dagger})$ by a gauge transformation — compare with (29) below.
The gauge-fixing condition (25) was not motivated by the divergence of moduli space metrics but instead arose from a study of supersymmetry transformations. Indeed, comments in the introduction of [11] suggest that the authors had not suspected that this gauge-fixing constraint would cure the metric divergence. Nevertheless, we will show below that this gauge-fixing constraint leads to a satisfying regularisation of the divergent metric. We will show in the next section that:
Theorem 3.
Let $(A,\Phi)$ be a hyperbolic monopole with half-integer mass and let $(a,\phi)$ be a solution of the linearised Bogomolny equation (20) that represents a tangent vector to the moduli space. Then there exists an unframed $\mathfrak{sl}(2,\mathbb{C})$ gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(a+d^{A}\chi,\phi+[\Phi,\chi])$ satisfies the gauge-fixing condition (25). This $\chi$ has the following properties:
  1. 1.
    Over $S^{2}_{\infty}$, $a_{-}$ is a section of $\Lambda^{1,0}$.
  2. 2.
    The $L^{2}$ norm (23) of $(a_{-},\phi_{-})$ is finite.
In order to apply this theorem to the framed moduli space, we need the gauge transformation $\chi$ to be framed. But it is easy to modify $\chi$ so that it is framed. It is straightforward to check that if $(a_{-},\phi_{-})$ solves the gauge-fixing condition (25) then so does $(a_{-}-c\,d^{A}\Phi,\phi_{-})$ for any constant $c$. So replacing $\chi$ in Theorem 3 with $\tilde{\chi}=\chi-c\Phi$ and setting
$c=\int_{H^{3}}\operatorname{Tr}(d^{A}\chi\wedge F^{A})/\int_{H^{3}}\operatorname{Tr}(d^{A}\Phi\wedge F^{A})=-\frac{1}{4\pi np}\int_{H^{3}}\operatorname{Tr}(d^{A}\chi\wedge F^{A})$ (26)
produces a framed gauge transformation that still solves (25). Hence from Theorem 3 we also obtain the existence part of the following:
Theorem 4.
Let $(A,\Phi)$ and $(a,\phi)$ be as in Theorem 3. Then there exists a unique framed gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(a+d^{A}\chi,\phi+[\Phi,\chi])$ satisfies the gauge-fixing condition (25). This $\chi$ has the two properties listed in Theorem 3.
We will show that $(a_{-}-c\,d^{A}\Phi,\phi_{-})$ has the two properties listed in Theorem 3 at the end of section 3. The uniqueness part of this Theorem will be proved in section 4. Turning to the unframed moduli space, we have that:
Corollary 5.
Let $(A,\Phi)$ and $(a,\phi)$ be as in Theorem 3. Then there exists a unique unframed gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(a+d^{A}\chi,\phi+[\Phi,\chi])$ satisfies the gauge-fixing condition (25) and $\int_{H^{3}}\operatorname{Tr}(a_{-}\wedge F^{A})=0$. This $\chi$ has the two properties listed in Theorem 3.
Proof.
We may write the unframed gauge transformation in the form $\chi=\chi_{0}-cd^{A}\Phi$, in which $\chi_{0}$ is a framed gauge transformation and $c\in\mathbb{C}$. The requirement $\int_{H^{3}}\operatorname{Tr}(a_{-}\wedge F^{A})=0$ is met if and only if $c=\int_{H^{3}}\operatorname{Tr}(a\wedge F^{A})/\int_{H^{3}}\operatorname{Tr}(d^{A}\Phi\wedge F^{A})$. Then Theorem 4 ensures existence and uniqueness of a suitable framed gauge transformation $\chi_{0}$. ∎
Theorem 4 yields two symmetric bilinear forms $g_{\pm}$ on the tangent bundle of the framed monopole moduli space. They are defined by
$g_{\pm}(X,X)=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(a_{\pm}\wedge\ast a_{\pm}+\phi_{\pm}\ast\phi_{\pm}),$ (27)
in which $(a_{-},\phi_{-})$ is the solution of the gauge fixing condition (25) determined in Theorem 4, and $(a_{+},\phi_{+})=(-a_{-}^{\dagger},-\phi_{-}^{\dagger})$. By construction, the bilinear forms $g_{\pm}$ are also related by conjugation: in other words,
$g_{-}(X,X)=\overline{g_{+}(X,X)}$ (28)
for any real tangent vector $X$.
We have been careful to avoid calling $g_{\pm}$ metrics because we do not know whether they always real-valued, and if so, whether they are positive definite. They are real if and only if $g_{+}=g_{-}$, so we can investigate reality by calculating their difference. To do this, it is convenient to write
$(a_{\pm},\phi_{\pm})=(a_{0}\pm id^{A}\lambda,\phi_{0}\pm i[\Phi,\lambda]),$ (29)
in which $a_{0},\phi,\lambda$ are $\mathfrak{su}(2)$-valued. The gauge-fixing condition (25) is then equivalent to
$\ast d^{A}\ast a_{0}+[\Phi,\phi_{0}]-2[\Phi,\lambda]$ $=0,$ (30)
$\ast d^{A}\ast d^{A}\lambda+[\Phi,[\Phi,\lambda]]+2\phi_{0}$ $=0.$ (31)
A short calculation shows that
$g_{+}(X,X)-g_{-}(X,X)$ $=-2i\int_{H^{3}}\operatorname{Tr}([\Phi,\lambda]\ast\phi_{0}+d^{A}\lambda\wedge\ast a_{0})$ (32)
$=2i\int_{H^{3}}\operatorname{Tr}(\lambda([\Phi,\ast\phi_{0}]+d^{A}\ast a_{0}))-d\operatorname{Tr}(\lambda\ast a_{0})$ (33)
$=4i\int_{H^{3}}\operatorname{Tr}\ast(\lambda[\Phi,\lambda])-2i\int_{S^{2}_{\infty}}\operatorname{Tr}(\lambda\ast a_{0})$ (34)
$=-2i\int_{S^{2}_{\infty}}\operatorname{Tr}(\lambda\ast a_{0}).$ (35)
Thus the imaginary part of $g_{+}$, which is proportional to $g_{+}-g_{-}$, reduces to an integral over the boundary of $H^{3}$. However, we have not been able to show in general that this boundary integral vanishes.
More can be said in the case of inversion-symmetric monopoles. Recall that inversion about a point in $H^{3}$ is the unique orientation-reversing isometry $I:H^{3}\to H^{3}$ whose fixed set is a point and which squares to the identity. A monopole is called inversion-symmetric if
$(A,\Phi)=(I^{\ast}A,-I^{\ast}\Phi)$ (36)
up to gauge equivalence (the minus sign in this equation ensures that the right-hand-side solves (3)). Now suppose that $(a_{\pm},\phi_{\pm})$ solve the gauge-fixing condition (25) and represent a chosen real tangent vector $X$ to the moduli space of inversion-symmetric monopoles. Then $(a_{\pm}^{I},\phi_{\pm}^{I}):=(I^{\ast}a_{\pm},-I^{\ast}\phi_{\pm})$ differ from $(a_{\pm},\phi_{\pm})$ by gauge transformations. It is straightforward to check that
$\ast d^{A}\ast a_{\pm}^{I}+[\Phi,\phi_{\pm}^{I}]\mp 2i\phi_{\pm}^{I}=0.$ (37)
By the uniqueness part of Theorem 4, this means that $(a_{\pm}^{I},\phi_{\pm}^{I})=(a_{\mp},\phi_{\mp})$, in other words,
$I^{\ast}a_{\pm}=a_{\mp}\quad\text{and}\quad I^{\ast}\phi_{\pm}=-\phi_{\mp}.$ (38)
Since the $L^{2}$ inner product (27) is invariant under inversions, this means that
$g_{-}(X,X)=g_{+}(X,X)$ (39)
for all real tangent vectors $X$ to the moduli space of inversion-symmetric monopoles. Combined with the relation (28), this shows that
Proposition 6.
$g_{+}=g_{-}$ defines a real symmetric bilinear form on the moduli space of inversion-symmetric hyperbolic monopoles.
A similar argument shows that the restriction of $g_{+}$ to the normal bundle of the moduli space of inversion-symmetric monopoles is also real. For normal vectors one has that $(a_{\pm}^{I},\phi_{\pm}^{I})$ is gauge equivalent to $(-a_{\pm},-\phi_{\pm})$ rather than $(a,\phi)$, and hence that $I^{\ast}a_{\pm}=-a_{\mp}$ and $I^{\ast}\phi_{\pm}=+\phi_{\mp}$, leading to the conclusion that $g_{+}(X,X)=g_{-}(X,X)$ for normal tangent vectors $X$. In the particular case of charge 1, it is known that every monopole is inversion-symmetric about some point and that the moduli-space of inversion-symmetric monopoles about a given point is 0-dimensional. Therefore every tangent vector is normal to some inversion-symmetric moduli space and thus $g_{\pm}$ is a real symmetric bilinear form on moduli space of 1-monopoles. We will calculate this metric explicitly in section 5 and confirm that it is positive-definite.
Turning our attention to the unframed moduli spaces, Corollary 5 allows us to define a pair of complex-conjugate symmetric bilinear forms via the integral (27). Equivalently, these bilinear forms could be obtained from those on the framed moduli spaces by the quotient construction. Again, we do not know whether these are real, but they are real when restricted to inversion-symmetric monopoles, and they are real and positive definite on the unframed moduli space of 1-monopoles as we show in Section 5.
3 Existence for gauge fixing condition
The purpose of this section is to prove Theorem 3, in other words, to show that the gauge-fixing condition (25) has a unique solution. We begin by reformulating it in a 4-dimensional way. Let $g_{H}=d\theta^{2}+g_{H^{3}}$ be the metric on $S^{1}\times H^{3}$ and let $\ast_{H}$ be the associated Hodge star. Let $B=\Phi\,d\theta+A$ be the circle-invariant connection on $S^{4}$ and let $b=\phi\,d\theta+a$ be the deformation of $B$ associated to $(a,\phi)$. Then, for the choice of orientation $d\theta\wedge\operatorname{Vol}_{H^{3}}$, (25) is equivalent to $\delta^{\pm}b=0$, where
$\delta^{\pm}:\Omega^{1}(\mathfrak{su}(2))\to\Omega^{0}(\mathfrak{su}(2)),\quad\delta^{\pm}b=e^{\mp 4i\theta}\ast_{H}d^{B}\ast_{H}e^{\pm 2i\theta}b$ (40)
Here the factor of $e^{\mp 4i\theta}$ is included for later notational convenience.
We have already seen in (7) that the metric $g_{H}$ on $S^{1}\times H^{3}$ is conformally equivalent to spherical metric $g_{S}=dy_{\mu}dy_{\mu}$ on $S^{4}\setminus S^{2}$. To begin analysing equation (40) we employ a second conformal equivalence. Introducing $r_{2}^{2}=y_{3}^{2}+y_{4}^{2}+y_{5}^{2}$, we note that
$\frac{1}{r_{2}^{2}}dy_{\mu}dy_{\mu}=\frac{1}{r_{2}^{2}}(dr_{2}^{2}+dy_{1}^{2}+dy_{2}^{2})+\frac{1}{r_{2}^{2}}(dy_{3}^{3}+dy_{4}^{2}+dy_{5}^{2})$ (41)
is the product metric on $H^{3}\times S^{2}$, with $S^{2}$ consisting of unit vectors $(y_{3},y_{4},y_{5})/r_{2}$. The $S^{4}$ constraint $y_{\mu}y_{\mu}=1$ is equivalent to $r_{2}^{2}+y_{1}^{2}+y_{2}^{2}=1$, which defines a copy of $H^{2}$ inside $H^{3}$. So the metric $g_{S}$ on $S^{4}\setminus S^{2}$ is conformal to the product metric $g_{K}$ on $H^{2}\times S^{2}$. It is convenient to work in the disc model of $H^{2}$, so we introduce a stereographic coordinate $w\in\mathbb{C}$ satisfying $|w|<1$, such that
$(r_{2},y_{1},y_{2})=\left(\frac{1-|w|^{2}}{1+|w|^{2}},\frac{\bar{w}+w}{1+|w|^{2}},\frac{i\bar{w}-i{w}}{1+|w|^{2}}\right).$ (42)
Then the argument of $w$ equals $\theta$ and the circle action fixes the $S^{2}$ given by $w=0$. The product metric on $H^{2}\times S^{2}$ obtained from (41) is
$g_{K}=\frac{2}{(1-|w|^{2})^{2}}(dw\,d{\bar{w}}+d{\bar{w}}d{w})+\frac{2}{(1+|z|^{2})^{2}}(dz\,d{\bar{z}}+d{\bar{z}}d{z}),$ (43)
with $z\in\mathbb{C}$ being a stereographic coordinate on $S^{2}$. Since $r_{2}=(1-|w|^{2})/(1+|w|^{2})$ and $r_{1}=2|w|/(1+|w|^{2})$, the metric $g_{H}$ of equation (7) and the spherical metric $g_{S}$ are related to $g_{K}$ by
$g_{H}=\left(\frac{1+|w|^{2}}{2|w|}\right)^{2}g_{S}=\left(\frac{1-|w|^{2}}{2|w|}\right)^{2}g_{K}.$ (44)
Moreover, the natural orientations of $S^{1}\times H^{3}$ and $H^{2}\times S^{2}$ agree.
Like $g_{S}$, the metric $g_{K}$ extends smoothly across the $S^{2}\subset S^{4}$ fixed by the circle action. We will use the metric $g_{K}$ to show that the gauge-fixing condition $\delta^{-}b=0$ admits a solution in a neighbourhood of this $S^{2}$, before showing that a solution exists on all of $S^{4}$. Solving the alternative condition $\delta^{+}b=0$ is equivalent, and we focus on $\delta^{-}$ for notational convenience. The advantage of working with $g_{K}$, rather than directly with the round metric $g_{S}$, is that $g_{K}$ is Kähler, with Kähler form
$\omega_{K}=\frac{2i}{(1-|w|^{2})^{2}}dw\wedge d{\bar{w}}+\frac{2i}{(1+|z|^{2})^{2}}dz\wedge d{\bar{z}}.$ (45)
The Bogomolony equation for $(A,\Phi)$ is equivalent to the anti-self-dual equations for $B$, which are conformally invariant and take the form
$(F^{B})^{0,2}=0=(F^{B})^{2,0},\quad\omega_{K}\wedge F^{B}=0$ (46)
in terms of the complex structure of $H^{2}\times S^{2}$. Their linearisation is
¯$\bar{\partial}^{B}b^{0,1}=0={\partial}^{B}b^{1,0},\quad\omega_{K}\wedge d^{B}b=0.$ (47)
Thus $B$ determines a holomorphic structure on the vector bundle $V$ over $H^{2}\times S^{2}$ and $b^{0,1}$ is a holomorphic section of $\operatorname{End}_{0}(V)\otimes\Lambda^{0,1}$, with the subscript 0 indicating that $b^{0,1}$ is traceless. This section in turn determines a cohomology class in $H^{0,1}(H^{2}\times S^{2},\operatorname{End}_{0}(V))$. In order to solve the gauge fixing condition it will be convenient to choose a gauge in which ${b}^{0,1}$ vanishes at $w=0$. To do so, we first need to determine the cohomology group $H^{0,1}(H^{2}\times S^{2},\operatorname{End}_{0}(V))$.
We can determine the cohomology of $\operatorname{End}_{0}(V)$ using a natural filtration of $V$, i.e. a holomorphic line bundle $L\subset V$. Following [1], this can be constructed either as the unique circle-invariant continuation of the bundle $L^{-}$ over the 2-sphere at $w=0$, or as the space of decaying solutions to a scattering equation along radial lines. We describe both constructions, beginning with the circle action.
The circle action on $H^{2}\times S^{2}$ is $(w,z)\mapsto(\lambda^{2}w,z)$, with $\lambda=e^{it/2}$. The lift of this action to $V$ fixes $B$ and so preserves the holomorphic structure. It therefore extends to a holomorphic action of the monoid $\mathbb{C}^{\ast}_{\leq 1}=\{\lambda\in\mathbb{C}^{\ast}\>:\>|\lambda|\leq 1\}$ (here the constraint on the modulus of $\lambda$ is imposed in order maintain the condition that $|w|<1$). From (10), the action on the subbundles $L^{\pm}$ over the fixed 2-sphere is given by $\lambda^{\mp 2p}$.
A line subbundle of $V$ corresponds to a section of $\mathbb{P}V$. The line subbundle $L^{-}$ over $\{0\}\times S^{2}$ determines a submanifold $C\subset\mathbb{P}V$ whose points are fixed by the action of $\mathbb{C}^{\ast}_{\leq 1}$. The tangent space of $\mathbb{P}V$ at a point in $C$ can be decomposed into weight spaces for the action, and the weights are $-4p$, 00 and $2$. The weight 00 corresponds to the tangent space of $C$, and the weight $-4p$ corresponds to the tangent space of the fibre of $\mathbb{P}V\to H^{2}\times S^{2}$. The stable submanifold of $C$, consisting of points in $\mathbb{P}V$ that converge to $C$ under the action of $\lambda$ as $\lambda\to 0$, has complex dimension 1. This submanifold is a section of $\mathbb{P}V$, because $\mathbb{C}^{\ast}_{\leq 1}$ acts freely and transitively on $H^{2}\times S^{2}$. It hence determines a line subbundle $L\subset V$. The degree of this line bundle $L$ equals the degree of its restriction to $S^{2}$, which is $-n$.
Now we present the alternative scattering description of $L$. From (43) and (44) we see that $s=-\ln|w|$ is an outward-pointing geodesic radial coordinate on $H^{3}$. Then $w=e^{-s+i\theta}$ is holomorphic coordinate on a surface of constant $z$. Holomorphic sections $\sigma$ on this surface satisfy $(\partial_{s}^{B}-i\partial_{\theta}^{B})\sigma=0$. In terms of the monopole, a holomorphic section which is invariant under the circle action corresponds to a solution of Hitchin’s scattering equation $(\frac{d}{ds}+A_{s}-i\Phi)\sigma=0$ along a geodesic ray parametrised by $s>0$. Solutions typically blow up like $e^{ps}$ as $s\to\infty$, but there is a one-dimensional space of solutions that decays like $e^{-ps}$. These solutions asymptote to $L^{-}$, the eigenspace of $\Phi$ with eigenvalue $ip$. The union of all these lines is the bundle $L$, and the Bogomolny equations ensure that it is a holomorphic subbundle.
The filtration $0\subset L\subset V$ determines a filtration of $\operatorname{End}_{0}(V)$
$0=F_{0}\subset F_{1}\subset F_{2}\subset F_{3}=\mathrm{End}_{0}(V),$ (48)
in which
$F_{1}$ =$=\{\chi\in\mathrm{End}_{0}(V)\>:\>\chi(V)\subset L,\,\chi(L)=0\}$ (49)
$F_{2}$ =$=\{\chi\in\mathrm{End}_{0}(V)\>:\>\chi(L)\subset L\}.$ (50)
Then $F_{i}$ has rank $i$ and $F_{1}$, $F_{2}/F_{1}$ and $F_{3}/F_{2}$ are line bundles of degree $-n$, 00 and $n$. They are isomorphic to the standard bundles $\mathcal{O}(-n)$, $\mathcal{O}$ and $\mathcal{O}(n)$ over $S^{2}$, pulled back to $H^{2}\times S^{2}$. These isomorphisms can be chosen so that $U(1)$ acts as $w\mapsto\lambda^{2}w$ on $H^{2}$ and by multiplication with $\lambda^{4p}$, $1$ and $\lambda^{-4p}$ on the fibres of $\mathcal{O}(-n)$, $\mathcal{O}$ and $\mathcal{O}(n)$ respectively.
With this notation in place, we are able to compute the Dolbeault cohomology of $\mathrm{End}_{0}(V)$. The result that we need is
Lemma 7.
The natural maps
$H^{0,0}(H^{2})\otimes H^{0,1}(S^{2},\mathcal{O}(-n))\to H^{0,1}(H^{2}\times S^{2},F_{1})\\ \to H^{0,1}(H^{2}\times S^{2},\operatorname{End}_{0}(V))$ (51)
are surjective.
Proof.
By the Künneth formula,
$H^{0,1}(H^{2}\times S^{2},F_{3}/F_{2})\cong H^{0,1}(H^{2}\times S^{2},\mathcal{O}(n))\\ \cong H^{0,0}(H^{2})\otimes H^{0,1}(S^{2},\mathcal{O}(n))\oplus H^{0,1}(H^{2})\otimes H^{0,0}(S^{2},\mathcal{O}(n)).$ (52)
This vanishes, because $H^{0,1}(S^{2},\mathcal{O}(n))=0$ for any $n>0$ and $H^{0,1}(H^{2})=0$. Therefore the map
$H^{0,1}(H^{2}\times S^{2},F_{2})\to H^{0,1}(H^{2}\times S^{2},F_{3})$ (53)
associated with the exact sequence $F_{2}\to F_{3}\to F_{3}/F_{2}$ is surjective. By a similar argument based on the fact that $H^{0,1}(S^{2},\mathcal{O})=0$, the map
$H^{0,1}(H^{2}\times S^{2},F_{1})\to H^{0,1}(H^{2}\times S^{2},F_{2})$ (54)
associated with the exact sequence $F_{1}\to F_{2}\to F_{2}/F_{1}$ is surjective. Applying the Künneth formula once more, we obtain a natural isomorphism
$H^{0,1}(H^{2}\times S^{2},F_{1})$ $\cong H^{0,1}(H^{2}\times S^{2},\mathcal{O}(-n))$ (55)
$=H^{0,0}(H^{2})\otimes H^{0,1}(S^{2},\mathcal{O}(-n))$ (56)
since $H^{0,1}(H^{2})=0$. ∎
The lemma implies that we can find a gauge transformation $\chi\in\Gamma(H^{2}\times S^{2},\operatorname{End}_{0}(V))$ such that ${b}^{0,1}-\bar{\partial}^{B}\chi$ equals a holomorphic function $f$ on $H^{2}$ times a section $\sigma$ of $\Lambda^{0,1}(S^{2})\otimes\mathcal{O}(-n)$. This can be done in such a way that ${b}-d^{B}\chi$ is still $U(1)$-invariant. Recalling that $U(1)$ acts on $\mathcal{O}(-n)$ as multiplication by $\lambda^{4p}$, this means that the function $f$ must satisfy $f(\lambda^{2}w)=\lambda^{4p}f(w)$. The only such holomorphic function is $f(w)=w^{2p}$. So we are able to choose a gauge in which
${b}^{0,1}=w^{2p}\sigma,\quad\sigma\in\Gamma(\Lambda^{0,1}(S^{2})\otimes\operatorname{End}_{0}(V)).$ (57)
Now let us investigate the implications of this choice for the gauge-fixing condition. The operator $\delta^{-}$ of (40) can be rewritten using the second conformal equivalence of (44) as follows:
$\delta^{-}=\left(\frac{2w}{1-|w|^{2}}\right)^{4}\ast_{K}d^{B}\ast_{K}\left(\frac{1-|w|^{2}}{2w}\right)^{2}.$ (58)
If ${b}$ satisfies (57) then this simplifies considerably. To see why, note that (58) implies that
$\delta^{-}b=\left(\frac{2w}{1-|w|^{2}}\right)^{2}\ast_{K}d^{B}\ast_{K}{b}-\left(\frac{2w}{1-|w|^{2}}\right)^{3}\left(\frac{dw}{w^{2}}+d\bar{w}\right)\lrcorner_{K}b.$ (59)
The contraction with $dw$ is 00 in the gauge (57) because $b_{\bar{w}}=0$. So the second term is $w^{3}$ times a smooth function. The first term can be simplified using the linearised anti-self-dual equations (47) and the fact that $d\omega_{K}=0$:
$\begin{split}d^{B}\ast_{K}{b}&=d^{B}(i\omega_{K}\wedge({b}^{0,1}-{b}^{1,0}))\\ &=d^{B}(\omega_{K}\wedge(2i{b}^{0,1}-i{b}))\\ &=2i\omega_{K}\wedge d^{B}{b}^{0,1}\\ &=2i\omega_{K}\wedge(2pw^{2p-1}dw\wedge\sigma+w^{2p}d^{B}\sigma)\\ &=2iw^{2p}\omega_{K}\wedge d^{B}\sigma.\end{split}$ (60)
Since $2p\geq 1$, the first term in (59) is also $w^{3}$ times a smooth function. Altogether, we have shown that
Lemma 8.
Suppose that the gauge is chosen so that ${b}$ satisfies (57) in a neighbourhood of $w=0$. Then
$\frac{1}{w^{3}}\delta^{-}b$ (61)
is a smooth function on $S^{4}$.
Having chosen this initial gauge, we now try to solve the gauge fixing condition 25. In other words, we seek a section $\chi$ of $\operatorname{End}_{0}(V)$ such that
$\delta^{-}(b+d^{B}\chi)=0,$ (62)
where $\delta^{-}$ is given in (40) and $b$ satisfies (57). We rewrite this using the following result
Lemma 9.
Let $\triangle_{S}^{B}=-\ast_{S}d^{B}\ast_{S}d^{B}$ denote the (positive) Hodge laplacian associated with the connection $B$ over $S^{4}$. For any section $\chi$ of $\operatorname{End}_{0}(V)$,
$\delta^{-}d^{B}\chi=-\left(\frac{2w}{1+|w|^{2}}\right)^{3}(\triangle^{B}_{S}+2)\left(\frac{1+|w|^{2}}{2w}\right)\chi.$ (63)
Proof.
The first conformal equivalence of (44) implies that
$\delta^{-}d^{B}\chi=\left(\frac{2w}{1+|w|^{2}}\right)^{4}\ast_{S}d^{B}\ast_{S}\left(\frac{1+|w|^{2}}{2w}\right)^{2}d^{B}\chi.$ (64)
The statement of the lemma is therefore equivalent to
$\ast_{S}d^{B}\ast_{S}\left(\frac{1+|w|^{2}}{2w}\right)^{2}d^{B}\chi=-\left(\frac{1+|w|^{2}}{2w}\right)(\triangle^{B}_{S}+2)\left(\frac{1+|w|^{2}}{2w}\right)\chi.$ (65)
We prove the identity (65) by expanding both sides. The left hand side is clearly equal to
$-\left(\frac{1+|w|^{2}}{2w}\right)^{2}\triangle_{S}^{B}\chi+2\left(\frac{1+|w|^{2}}{2w}\right)\ast_{S}\left(d\left(\frac{1+|w|^{2}}{2w}\right)\wedge\ast_{S}\,d^{B}\chi\right).$ (66)
The right hand side equals
$-\left(\frac{1+|w|^{2}}{2w}\right)^{2}\triangle_{S}^{B}\chi+2\left(\frac{1+|w|^{2}}{2w}\right)\ast_{S}\left(d\left(\frac{1+|w|^{2}}{2w}\right)\wedge\ast_{S}\,d^{B}\chi\right)\\ -\left(\frac{1+|w|^{2}}{2w}\right)\chi(\triangle_{S}+2)\left(\frac{1+|w|^{2}}{2w}\right).$ (67)
A direct calculation shows that
$\triangle_{S}\left(\frac{1+|w|^{2}}{2w}\right)=-2\left(\frac{1+|w|^{2}}{2w}\right),$ (68)
so the two sides of the identity agree. ∎
Lemma 9 reduces the gauge fixing condition (62) to
$(\triangle_{S}^{B}+2)\left(\frac{1+|w|^{2}}{2w}\right)\chi=\left(\frac{1+|w|^{2}}{2w}\right)^{3}\delta^{-}b.$ (69)
Lemma 8 shows that $b$ can be chosen so that the right hand side is a smooth function near $S^{2}\subset S^{4}$. Standard theory guarantees that (69) admits a unique smooth solution. In more detail, the operator $\triangle_{S}^{B}+2$ is a self-adjoint second order elliptic operator. Its kernel is 0, because
$(\triangle^{B}_{S}+2)f=0\quad\implies\quad\int_{S^{4}}\left(|d^{B}f|^{2}+2|f|^{2}\right)\mathrm{Vol}_{S}=0\quad\implies\quad f=0.$ (70)
By the Fredholm alternative, (69) admits a unique solution $\chi$. By uniqueness, this solution is circle-invariant. So $b_{-}=b+d^{B}\chi$ solves the gauge-fixing condition (62) and is circle-invariant. It gives a solution $(a_{-},\phi_{-})$ of the gauge-fixing condition (25) with a minus sign. This proves the existence part of Theorem 3.
Next we investigate the behaviour of the solution near $S^{2}$. From (57) we see that $b^{0,1}$ is equal to $w$ times a smooth section near $w=0$. From (69) we see that $\chi$ is also equal to $w$ times a smooth section near $w=0$. Since $w$ is a holomorphic coordinate, the $(0,1)$ part of $d^{B}\chi$ is also equal to $w$ times a smooth section, and so the $(0,1)$ part of $b_{-}=b+d^{B}\chi$ equals $w$ times a smooth section. This proves part of the following lemma:
Lemma 10.
The solution $(a_{-},\phi_{-})$ of (25) in Theorems 3 and 4 is such that the (0,1)-part of $b_{-}=a_{-}+\phi_{-}\,d\theta$ is equal to $w$ times a smooth section near $S^{2}\subset S^{4}$.
Proof.
The statement is already proved in the case of Theorem 3. For Theorem 4 we in addition need to show that the (0,1)-part of $d^{A}\Phi$ is $w$ times a smooth section. The connection $B=A+\Phi\,d\theta$ has curvature $F^{B}=F^{A}+d^{A}\Phi\wedge d\theta$, so
$d^{A}\Phi=-\frac{\partial}{\partial\theta}\lrcorner F^{B}=-iw\frac{\partial}{\partial w}\lrcorner F^{B}+i\bar{w}{\frac{\partial}{\partial\bar{w}}}\lrcorner F^{B}.$ (71)
The anti-self-dual equations (46) imply that $F^{B}$ is of type (1,1), so the (0,1)-part of $d^{A}\Phi$ is $-i{w}{\frac{\partial}{\partial{w}}}\lrcorner F^{B}$. Thus $d^{A}\Phi$ is $w$ times a smooth section. ∎
Now we show that the metric integral (23) is finite for the solution $(a_{-},\phi_{-})$ of the gauge fixing condition (25). This integral is proportional to
$-\int_{S^{4}\setminus S^{2}}\operatorname{Tr}(b_{-}\wedge\ast_{H}b_{-})=-\int_{S^{4}\setminus S^{2}}\left(\frac{1+|w|^{2}}{2|w|}\right)^{2}\operatorname{Tr}(b_{-}\wedge\ast_{S}b_{-}).$ (72)
The integrand is continuous away from $w=0$, so the integral will be finite as long as the integral in a neighbourhood of 00 is finite. The integral on a neighbourhood $0<|w|<\epsilon$ can be written using the Kähler metric (43) as follows:
$-\int_{0<|w|<\epsilon}\operatorname{Tr}(b_{-}\wedge\ast_{H}b_{-})=-2\int_{0<|w|<\epsilon}\left(\frac{1-|w|^{2}}{2|w|}\right)^{2}\operatorname{Tr}(b_{-}^{1,0}\wedge\ast_{K}b_{-}^{0,1}).$ (73)
[…]
4 Killing spinors and pluricomplex structures
The gauge fixing condition (25) was motivated by a connection with Killing spinors [11]. We will present a precise formulation of this correspondence below, and show that it induces geometric structures previously studied by Nash and Bielawski–Schwachhofer. We then show that these geometric structures are compatible with the symmetric bilinear forms $g_{\pm}$.
A Killing spinor on a riemannian manifold $M$ with a spin structure is a section $\psi$ of the spinor bundle $S$ that solves
$\nabla_{X}\psi=\lambda X\cdot\psi$ (74)
for all tangent vectors $X$. Here $\lambda\in\mathbb{C}$ is known as the Killing constant of $\psi$, and the dot denotes Clifford multiplication.
To relate this to monopoles, let us suppose that $M$ is a three-dimensional. Then the Clifford algebra admits two inequivalent representations; we choose the one in which the volume form acts as the identity, and this choice determines the spinor bundle $S$. We recall that $S$ is a complex rank 2 bundle admitting a non-degenerate skew-symmetric pairing (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). Such a pairing induces an isomorphism $S\cong S^{\ast}$, written as $s\mapsto s^{\ast}$, such that the pairing is $(s,t)=s^{\ast}(t)$. In a suitable local frame the map $s\mapsto s^{\ast}$ takes the form
$\begin{pmatrix}s_{1}\\ s_{2}\end{pmatrix}^{\ast}=\begin{pmatrix}-s_{2}&s_{1}\end{pmatrix},$ (75)
corresponding to the skew-symmetric pairing
$(s,t)=\begin{pmatrix}s_{1}&s_{2}\end{pmatrix}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\begin{pmatrix}t_{1}\\ t_{2}\end{pmatrix}=s_{1}t_{2}-s_{2}t_{1}.$ (76)
The skew-symmetric pairing is compatible with Clifford multiplication in the sense that $(s,v\cdot t)=-(v\cdot s,t)$ and $(v\cdot s)^{\ast}=-s^{\ast}\cdot v$ for all $v\in\Lambda^{1}$.
Clifford multiplication induces an isomorphism
$\operatorname{cl}_{H}:(\Lambda^{0}\oplus\Lambda^{1})\otimes\mathbb{C}\to\operatorname{End}(S)\cong S\otimes S^{\ast}\cong S\otimes S$ (77)
such that $(a+\phi)\cdot\psi=\operatorname{cl}_{H}(a,\phi)\psi$ for all $a\in\Lambda^{1},\phi\in\Lambda^{0},\psi\in S$. Under this isomorphism $\Lambda^{1}$ corresponds to the traceless endomorphisms $\operatorname{End}_{0}(S)$ and $\Lambda^{0}$ corresponds to the identity component.
Three-dimensional hyperbolic space $H^{3}$ admits Killing spinors with Killing constant ±$\pm\frac{i}{2}$. The space $K^{\pm}$ of Killing spinors with Killing constant ±$\pm\frac{i}{2}$ is two-dimensional. Since the spinor bundle $S$ is two-dimensional and Killing spinors are determined by their value at a point, Killing spinors trivialise $S$. We thus have the following natural isomorphisms of vector bundles, for $p\in H^{3}$, $\psi^{\pm}\in K^{\pm}$,
$K^{\pm}\times H^{3}\cong S,\quad(\psi^{\pm},p)\mapsto\psi^{\pm}(p).$ (78)
It follows that
$\Gamma(\Lambda^{0}\oplus\Lambda^{1})\otimes\mathbb{C}\cong\Gamma(S)\otimes K^{\pm}.$ (79)
More explicitly, this means that given a pair of independent Killing spinors $\psi_{1},\psi_{2}$ in $K^{-}$, any section $(a,\phi)$ of $(\Lambda^{1}\oplus\Lambda^{0})\otimes\mathbb{C}$ can be written uniquely as in the form $a+\phi=\nu_{1}\otimes\psi_{1}^{\ast}+\nu_{2}\otimes\psi_{2}^{\ast}$ for a pair of sections $\nu_{1},\nu_{2}$ of $S$ (and similarly for $K^{+}$).
Proposition 11.
Let $(A,\Phi)$ be a hyperbolic monopole and let $D^{A}:\Gamma(S\otimes\operatorname{End}(V))\to\Gamma(S\otimes\operatorname{End}(V))$ be the Dirac operator associated to $A$. Under the the isomorphism (79), we have an identity
$\left(-D^{A}+\operatorname{ad}(\Phi)\pm\frac{i}{2}\right)\otimes\mathrm{id}_{K^{\pm}}=\begin{pmatrix}*d^{A}+\operatorname{ad}(\Phi)&-d^{A}\\ *d^{A}*&\operatorname{ad}(\Phi)\pm 2i\end{pmatrix},$ (80)
in which the operator on the right is written with respect to the splitting $\Lambda^{1}\oplus\Lambda^{0}$.
Proof.
Choose an orthonormal frame $e^{i}$ for the cotangent bundle, and choose a basis $\psi_{1},\psi_{2}$ for $K^{\pm}$. To simplify notation, write $\lambda=\pm\frac{i}{2}$. Then for any sections $\nu_{1},\nu_{2}$ of $S\otimes\operatorname{End}(V)$,
$e^{i}\cdot\nabla^{A}_{i}(\nu_{\alpha}\otimes\psi^{\ast}_{\alpha})$ $=e^{i}\cdot(\nabla^{A}_{i}\nu_{\alpha})\otimes\psi^{\ast}_{\alpha}+\lambda e^{i}\cdot\nu_{\alpha}\otimes(e^{i}\cdot\psi_{\alpha})^{\ast}$ (81)
$=(D^{A}\nu_{\alpha})\otimes\psi_{\alpha}^{\ast}-\lambda e^{i}\cdot\nu_{\alpha}\otimes\psi_{\alpha}^{\ast}\cdot e^{i}$ (82)
$=(D^{A}\nu_{\alpha}-\lambda\nu_{\alpha})\otimes\psi_{\alpha}^{\ast}+2\lambda\operatorname{Tr}_{S}(\nu_{\alpha}\otimes\psi_{\alpha}^{\ast}),$ (83)
in which we used the identity $e^{i}\cdot x\cdot e^{i}=x-2\operatorname{Tr}_{S}(x)$ for $x\in\mathrm{End}(S)$. Writing $\nu_{\alpha}\otimes\psi^{\ast}_{\alpha}=a+\phi$ for sections $a,\phi$ of $\Lambda^{1}\otimes\operatorname{End}(V)$ and $\Lambda^{0}\otimes\operatorname{End}(V)$, it follows that
$(-D^{A}\nu_{\alpha}+[\Phi,\nu_{\alpha}]+\lambda\nu_{\alpha})\otimes\psi_{\alpha}^{\ast}=-e^{i}\cdot\nabla^{A}_{i}(a+\phi)+[\Phi,a+\phi]+4\lambda\phi,$ (84)
having used $\operatorname{Tr}_{S}(\nu_{\alpha}\otimes\psi^{\ast}_{\alpha})=\operatorname{Tr}_{S}(\phi I_{2}+a^{i}e_{i}\cdot)=2\phi$. By direct calculation,
$-e^{i}\cdot\nabla^{A}_{i}({\phi}+a)=-e^{i}\nabla_{i}^{A}{\phi}-e^{i}\wedge\nabla_{i}^{A}a+e^{i}\lrcorner\nabla_{i}^{A}{a}=-d^{A}{\phi}-d^{A}{a}+*d^{A}*{a}.$ (85)
Since we have chosen the Clifford algebra representation in which the volume form $\mathrm{Vol}$ acts as multiplication by 1,
$-d^{A}{a}=-\mathrm{Vol}\cdot d^{A}{a}=*d^{A}{a}.$ (86)
The result follows from equations (84), (85) and (86). ∎
Proposition (11) gives a relationship between eigenspinors $\nu$ solving
$D^{A}\nu-[\Phi,\nu]=\pm\frac{i}{2}\nu$ (87)
and solutions $(a,\phi)$ of (20), (25) that describe tangent vectors to the moduli space. To make this relationship precise, we need to impose boundary conditions on $\nu$. Just as we did with tangent vectors $(a,\phi)$, we will impose these boundary conditions by lifting to $S^{4}$.
Recall that the spinor bundle of $S^{1}\times H^{3}$ takes the form $S_{+}\oplus S_{-}$. Self-dual two-forms act trivially on $S_{-}$, while anti-self-dual 2-forms act trivially on $S_{+}$. Both $S_{+}$ and $S_{-}$ are isomorphic to the pull-back of the spinor bundle $S$ of $H^{3}$. The Dirac operator associated to the connection $\Phi\,d\theta+A$ splits as the sum of an operator $D^{B}_{H}:\Gamma(S^{1}\times H^{3},S_{-}\otimes\operatorname{End}_{0}(V))\to\Gamma(S^{1}\times H^{3},S_{+}\otimes\operatorname{End}_{0}(V))$ and its adjoint. We focus on $D^{B}_{H}$, which takes the form
$D^{B}_{H}=D^{A}-\frac{\partial}{\partial\theta}-\operatorname{ad}(\Phi).$ (88)
So the eigenspinor equation (87) is equivalent to
$D^{B}_{H}e^{\pm\frac{i\theta}{2}}\nu=0.$ (89)
Recall that the metrics $g_{S}$ of $S^{4}$ and $g_{H}$ of $S^{1}\times H^{3}$ are conformally equivalent (44). It follows that their Dirac operators are related by
$D^{B}_{H}=\left(\frac{2|w|}{1+|w|^{2}}\right)^{\frac{5}{2}}D^{B}_{S}\left(\frac{2|w|}{1+|w|^{2}}\right)^{-\frac{3}{2}}.$ (90)
For $x\in M^{f}_{n,p}$ we define $E_{x}^{\pm}$ to be the space of solutions of (87) such that
$\tilde{\nu}:=\left(\frac{2|w|}{1+|w|^{2}}\right)^{-\frac{3}{2}}\nu e^{\pm\frac{i\theta}{2}}$ (91)
extends smoothly to $S^{4}$. Then $E_{x}$ is naturally isomorphic to the kernel of $D^{B}_{S}:\Gamma(S^{4},S_{-}\otimes\operatorname{End}_{0}(V))\to\Gamma(S^{4},S_{+}\otimes\operatorname{End}_{0}(V))$.
By construction, the spinor $\tilde{\nu}$ transforms with weight ±$\pm 1/2$ under the circle action. This fact allows us to determine the dimension of $E_{x}^{\pm}$ using the equivariant index theorem:
Proposition 12.
The complex vector spaces $E_{x}^{\pm}$ have dimension $2n$.
Proof.
Let $U(1)$ be the double-cover of the circle group acting on $S^{4}$. For $\gamma\in U(1)$, we define
$\text{ind}_{U(1)}(\gamma,D^{B}_{S})=\operatorname{Tr}(\gamma,\ker D^{B}_{S})-\operatorname{Tr}(\gamma,\operatorname{coker}D^{B}_{S}).$ (92)
The equivariant index theorem says that
$\text{ind}_{U(1)}(\gamma,D^{B}_{S})=\frac{1}{2\pi i}\int_{S^{2}}\frac{\hat{A}(S^{2})\ \text{ch}_{U(1)}(\gamma,\operatorname{End}_{0}(V))}{\text{ch}_{U(1)}(\gamma,S^{+}(NS^{2}))-\text{ch}_{U(1)}(\gamma,S^{-}(NS^{2}))}.$ (93)
Here $S(NS^{2})=S^{+}(NS^{2})\oplus S^{-}(NS^{2})$ is the spinor bundle associated with the normal bundle of $S^{2}\subset S^{4}$. The equivariant Chern character of a bundle $W\to S^{2}$ is simply
$\text{ch}_{U(1)}(\gamma,W):=\operatorname{Tr}_{E}(\gamma\cdot\exp(-F)),$ (94)
in which $F$ is the curvature of a connection on $W$. The $\hat{A}$ genus of $S^{2}$ is given in terms of its curvature tensor $R$ simply by
$\hat{A}(S^{2})=\left(\det\frac{R/2}{\sinh R/2}\right)^{\frac{1}{2}}=1-\frac{1}{48}\operatorname{Tr}_{TS^{2}}(R^{2})+\ldots=1.$ (95)
To calculate the equivariant Chern character, recall that $V\to S^{2}$ is isomorphic to $L^{+}\oplus L^{-}$ with $L^{-}=(L^{+})^{\ast}$, and so the bundle $\operatorname{End}_{0}(V)\to S^{2}$ is isomorphic to $(L^{+})^{\otimes 2}\oplus\mathbb{C}\oplus(L^{-})^{\otimes 2}$. The degrees of these bundles are $2n,0,-2n$ and their weights under the $U(1)$ action are $-4p,0,4p$. We choose a connection on $L^{+}$ with curvature $-i\omega/2$, where $\omega$ is the area form on $S^{2}$ whose integral is $4\pi$. Then
$\text{ch}_{U(1)}(\gamma,\operatorname{End}_{0}(V))=\gamma^{-4p}+1+\gamma^{4p}+in\omega(\gamma^{4p}-\gamma^{-4p}).$ (96)
The normal bundle $NS^{2}$ is topologically trivial, and so $S^{\pm}(NS^{2})$ are also trivial. The action of the circle group on $NS^{2}$ has weights ±$\pm 1$, and the action of its double cover $U(1)$ has weights ±$\pm 2$, so the action of $U(1)$ on the positive and negative spinor bundles has weights ±$\pm 1$. Thus:
$\text{ch}_{U(1)}(\gamma,S^{\pm}(NS^{2}))=\gamma^{\pm 1}.$ (97)
Altogether, we have that
$\text{ind}_{U(1)}(\gamma,D^{B}_{S})=\frac{n}{2\pi}\int_{S^{2}}\frac{(\gamma^{4p}-\gamma^{-4p})}{\gamma-\gamma^{-1}}\omega=2n\sum_{j=0}^{4p-1}\gamma^{2j+1-4p}.$ (98)
The spaces $E_{x}^{\pm}$ are the weight spaces in $\ker D^{B}_{S}$ with weight ±$\pm 1$ under the double cover $U(1)$ of the circle group. Since the connection $B$ is anti-self-dual, a well-known argument shows that the cokernel of $D^{B}_{S}$ is trivial [3]. Therefore the equivariant index (92) is equal to the character of the representation $\ker(D^{B}_{S})$ of $U(1)$. The dimensions of the weight spaces $E^{\pm}_{x}$ are the coefficients $\gamma^{\pm 1}$ in this polynomial, and from the calculation (98) they are both equal to $2n$ as claimed. ∎
Returning to the discussion of boundary conditions, let $\nu\in E_{x}^{-}$ and let $\psi\in K^{-}$. From these we construct a solution $(a_{-},\phi_{-})$ on $S^{1}\times H^{3}$ of (20) and (25) via the isomorphism (77). We wish to show that $b_{-}:=\phi_{-}\,d\theta+a_{-}$ extends over $S^{2}$ to give a smooth section of $\Lambda^{1}S^{4}\otimes\operatorname{End}_{0}(V)$. Using the Clifford isomorphism
$\operatorname{cl}_{S}:\Lambda^{1}S^{4}\to\operatorname{Hom}(S_{+},S_{-})\cong S_{-}\otimes S_{+}^{\ast}\cong S_{-}\otimes S_{+}$ (99)
induced by the metric $g_{S}$ on $S^{4}$, $b_{-}$ can be thought of as a section of $S_{-}\otimes S_{+}\otimes\operatorname{End}_{0}(V)$, defined away from $S^{2}\subset S^{4}$. The Clifford isomorphisms (77), (99) are related by
$\operatorname{cl}_{S}=\frac{1+|w|^{2}}{2|w|}\operatorname{cl}_{H},$ (100)
in which we have implicitly identified $\Lambda^{1}(S^{4}\setminus S^{2})$ with $\Lambda^{1}(H^{3})\oplus\Lambda^{0}(H^{3})$ and identified $S$ with $S_{+}$ and $S_{-}$. So $\operatorname{cl}_{H}(b_{-})$ is given by
$\left(\frac{1+|w|^{2}}{2|w|}\right)\nu\otimes\psi,$ (101)
in which $\nu$ and $\psi$ are interpreted as sections of $S_{-}\otimes\operatorname{End}_{0}(V)$ and $S_{+}$. The section (101) can be written $\tilde{\nu}\otimes\tilde{\psi}$, with $\tilde{\nu}$ given in (91) and
$\tilde{\psi}:=\left(\frac{2|w|}{1+|w|^{2}}\right)^{\frac{1}{2}}e^{i\frac{\theta}{2}}\psi.$ (102)
Since $\tilde{\nu}$ is a smooth section over $S^{4}$, $b$ will extend smoothly over $S^{4}$ if $\tilde{\psi}$ extends to a smooth section of $S_{+}\to S^{4}$. The simplest way to show that $\tilde{\psi}$ extends smoothly is by explicitly solving the Killing spinor equation.
We do so in the upper half-space model of $H^{3}$. Define coordinates $\rho>0,x,y$ by stereographic projection:
$(y_{1},y_{2},y_{3},y_{4},y_{5})=\frac{1}{1+\rho^{2}+x^{2}+y^{2}}(2\rho\cos\theta,2\rho\sin\theta,2x,2y,1-\rho^{2}-x^{2}-y^{2}).$ (103)
Then the conformal factor in (44) is $2|w|/(1+|w|^{2})=r_{1}=2\rho/(1+\rho^{2}+x^{2}+y^{2})$. The hyperbolic metric and volume form are
$g=\frac{1}{\rho^{2}}(d\rho^{2}+dx^{2}+dy^{2}),\quad\mathrm{Vol}=-\frac{1}{\rho^{3}}d\rho\wedge dx\wedge dy.$ (104)
We choose a frame for the spinor bundle so that the Clifford isomorphism (77) is given by Pauli matrices:
$\operatorname{cl}_{H}:\frac{d\rho}{\rho}\mapsto-i\sigma_{3},\quad\frac{dx}{\rho}\mapsto-i\sigma_{1},\quad\frac{dy}{\rho}\mapsto-i\sigma_{2}.$ (105)
This ensures that $\mathrm{Vol}\mapsto 1$. Then the Levi-Civita connection on the spinor bundle is
$\nabla=d+\frac{dx}{2\rho}i\sigma_{2}-\frac{dy}{2\rho}i\sigma_{1}.$ (106)
A quick proof that this is the Levi-Civita connection is as follows: we compute
$[\nabla,i\sigma_{1}]=\frac{dx}{\rho}i\sigma_{3},\quad[\nabla,i\sigma_{2}]=\frac{dy}{\rho}i\sigma_{3},\quad[\nabla,i\sigma_{3}]=-\frac{dx}{\rho}i\sigma_{1}-\frac{dy}{\rho}i\sigma_{2}.$ (107)
So
$\nabla\frac{dx}{\rho}=\frac{dx}{\rho}\otimes\frac{d\rho}{\rho},\quad\nabla\frac{dy}{\rho}=\frac{dy}{\rho}\otimes\frac{d\rho}{\rho},\quad\nabla\frac{d\rho}{\rho}=-\frac{dx}{\rho}\otimes\frac{dx}{\rho}-\frac{dy}{\rho}\otimes\frac{dy}{\rho}.$ (108)
It is straightforward to check that this connection on $\Lambda^{1}(\mathbb{H}^{3})$ is orthogonal. It is torsion-free because $\nabla\wedge v=dv$ for $v=dx/\rho,dy/\rho,d\rho/\rho$. It is straightforward to check that
$\psi_{1}=\rho^{-\frac{1}{2}}\begin{pmatrix}1\\ 0\end{pmatrix},\quad\psi_{2}=\rho^{-\frac{1}{2}}\begin{pmatrix}-x+iy\\ \rho\end{pmatrix}$ (109)
are Killing spinors with Killing constant $-\frac{\mathrm{i}}{2}$ which form a basis for $K^{-}$.
Now we aim to write $\psi_{\alpha}$ as a spinors on $S^{4}\setminus S^{2}$. It is natural to identify $S^{1}\times H^{3}$ with $\mathbb{R}^{4}\setminus\mathbb{R}^{2}$ using polar coordinates $(\rho\cos\theta,\rho\sin\theta,x,y)$. We choose a frame for the spinor bundle of $S^{4}$ over $\mathbb{R}^{4}$ such that the Clifford isomorphism (99) is given by
$\operatorname{cl}_{S}:\frac{2d(\rho\cos\theta)}{1+\rho^{2}+x^{2}+y^{2}}$ $\mapsto-i\sigma_{3},$ $\frac{2d(\rho\sin\theta)}{1+\rho^{2}+x^{2}+y^{2}}$ $\mapsto I_{2},$ (110)
$\frac{2dx}{1+\rho^{2}+x^{2}+y^{2}}$ $\mapsto-i\sigma_{1},$ $\frac{2dy}{1+\rho^{2}+x^{2}+y^{2}}$ $\mapsto-i\sigma_{2}.$
Comparing (110) with (105) and the identification $\operatorname{cl}_{H}(d\theta)=\operatorname{cl}_{H}(1)=I_{2}$, we see that
$\operatorname{cl}_{S}(\cdot)=\frac{2\rho}{1+\rho^{2}+x^{2}+y^{2}}\exp(\tfrac{i\theta}{2}\sigma_{3})\operatorname{cl}_{H}(\cdot)\exp(\tfrac{i\theta}{2}\sigma_{3})$ (111)
in our chosen representations. By comparing this with (100) we can see how our frames for $S$ and $S_{\pm}$ are related. Thus the Killing spinors $\psi_{\alpha}$ written in the frame (105) for $S$ correspond to $\exp(-\frac{i\theta}{2}\sigma_{3})\psi_{\alpha}$ in the frame (110) for $S_{+}$. Then from (102) and (109) we find
$\tilde{\psi_{1}}=\left(\frac{4}{1+\rho^{2}+x^{2}+y^{2}}\right)^{\frac{1}{4}}\begin{pmatrix}1\\ 0\end{pmatrix},\,\tilde{\psi_{2}}=\left(\frac{4}{1+\rho^{2}+x^{2}+y^{2}}\right)^{\frac{1}{4}}\begin{pmatrix}-x+iy\\ \rho e^{i\theta}\end{pmatrix}.$ (112)
These are smooth functions on $\mathbb{R}^{4}$ and they are written in a frame that extends over $\mathbb{R}^{2}\subset\mathbb{R}^{4}$, so we conclude that the spinors $\tilde{\psi}_{\alpha}$ extend smoothly over $\mathbb{R}^{2}\subset\mathbb{R}^{4}$. Repeating the calculation in a different coordinate patch shows that the Killing spinors extend smoothly over $S^{2}\subset S^{4}$.
So if $\nu\in E_{x}^{-}$ and $\psi\in K^{-}$ then $(a_{-},\phi_{-}):=\operatorname{cl}_{H}^{-1}(\nu\otimes\psi^{\ast})$ is a solution of the linearised Bogomolny equations (20) and the gauge fixing condition (25) that extends smoothly over $S^{2}\subset S^{4}$. Similarly, $E^{+}\otimes K^{+}$ gives solutions of the “+” gauge fixing condition. We thus have natural maps to the complexified tangent space:
$E_{x}^{\pm}\otimes K^{\pm}\to T_{x}^{\mathbb{C}}M^{f}_{n,p}.$ (113)
We now wish to show that these maps are isomorphisms.
We begin by considering surjectivity. By Theorem 4, any tangent vector can be represented by a pair $(a_{-},\phi_{-})$ solving equations (20) and (25). We wish to write this pair in the form
$\operatorname{cl}_{H}(a_{-},\phi_{-})=\nu_{1}\otimes\psi_{1}^{\ast}+\nu_{2}\otimes\psi_{2}^{\ast}$ (114)
for some eigenspinors $\nu_{i}$ and Killing spinors $\psi_{i}$. We begin by choosing a pair of linearly independent Killing spinors. The Killing spinor equation implies that the function $(\psi_{1},\psi_{2})$ is constant; without loss of generality we can assume that this constant is 1 (it is certainly 1 for the choice made in (109)). Then $\psi_{2}\otimes\psi_{1}^{\ast}-\psi_{1}\otimes\psi_{2}^{\ast}\in S\otimes S^{\ast}$ corresponds to the identity map acting on $S$. So (114) holds with
$\nu_{\alpha}=\epsilon_{\alpha\beta}\operatorname{cl}_{H}(a_{-},\phi_{-})\psi_{\beta},$ (115)
where $\epsilon_{\alpha\beta}$ is totally antisymmetric and $\epsilon_{12}=1$. By Proposition 11, $\nu_{\alpha}$ solve the eigenspinor equation (87), so to confirm that they belong to $E_{x}^{-}$ we just need to check that $\tilde{\nu}_{\alpha}$ extend smoothly across $S^{2}\subset S^{4}$.
From equations (91), (100) and (102),
$\tilde{\nu}_{\alpha}$ $=\epsilon_{\alpha\beta}\left(\frac{2|w|}{1+|w|^{2}}\right)^{-\frac{3}{2}}e^{-\frac{i\theta}{2}}\frac{2|w|}{1+|w|^{2}}\operatorname{cl}_{S}(b_{-})\left(\frac{2|w|}{1+|w|^{2}}\right)^{-\frac{1}{2}}e^{-\frac{i\theta}{2}}\tilde{\psi}_{\beta}$ (116)
$=\frac{1+|w|^{2}}{w}\epsilon_{\alpha\beta}\operatorname{cl}_{S}(b_{-})\tilde{\psi}_{\beta}.$ (117)
In order to show that $\tilde{\nu}_{\alpha}$ extends smoothly, we need to show that $\operatorname{cl}_{S}(b_{-})\tilde{\psi}_{\beta}=O(w)$, in which the notation $O(w)$ means $w$ times a function that extends smoothly over $\mathbb{R}^{4}\setminus\mathbb{R}^{2}$. Recall from Lemma 10 that $b_{-}^{0,1}=O(w)$. Turning to $b_{-}^{1,0}$, we deduce from $w/(1+|w|^{2})=\rho e^{i\theta}/(1+\rho^{2}+x^{2}+y^{2})$ that $dw=d(\rho e^{i\theta})+O(w)$. So (110) gives
$\operatorname{cl}_{S}:\frac{dw}{1+\rho^{2}+x^{2}+y^{2}}\mapsto\begin{pmatrix}O(w)&O(w)\\ O(w)&i\end{pmatrix},\,\frac{dz}{1+\rho^{2}+x^{2}+y^{2}}\mapsto\begin{pmatrix}0&i\\ 0&0\end{pmatrix}.$ (118)
From this and equation (102) we deduce that
$\operatorname{cl}_{S}(b_{-})=\operatorname{cl}_{S}(b_{-}^{1,0})+O(w)=\begin{pmatrix}O(w)&O(1)\\ O(w)&O(1)\end{pmatrix},\quad\tilde{\psi}_{\beta}=\begin{pmatrix}O(1)\\ O(w)\end{pmatrix}.$ (119)
So $\operatorname{cl}_{S}(b_{-})\tilde{\psi}_{\beta}$ equals a smooth section times $w$, and hence $\tilde{\nu}_{\alpha}$ given in (117) extends smoothly across $\mathbb{R}^{2}\subset\mathbb{R}^{4}$. Repeating the calculation in a different coordinate patch shows that $\tilde{\nu}_{\alpha}$ extends smoothly over $S^{2}\subset S^{4}$. So $\nu_{\alpha}\in E_{x}^{-}$, and we have shown that (113) is surjective. Since the domain and target both have dimension $4n$, we conclude that
Theorem 13.
The maps given in (113) are isomorphisms.
A similar argument allows us to show that the framed gauge transformation that solves the gauge-fixing condition in Theorem 4 is unique. Suppose to the contrary: then there is a non-zero framed gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(d^{A}\chi,0)$ solves (20) and (25). This can be written in the form (114) with eigenspinors $\nu_{\alpha}\in E_{x}$ given in (115). Since the map (113) is injective and $(d^{A}\chi,0)$ represents the zero-vector in the quotient (22), it must be that $\nu_{\alpha}=0$. Then, from (114), $d^{A}\chi=0$. Since monopoles are irreducible this in turn implies that $\chi=0$, which contradicts our assumption.
[…]
$s=\begin{pmatrix}s_{1}\\ s_{2}\end{pmatrix}\mapsto\begin{pmatrix}\bar{s_{2}}\\ -\bar{s_{1}}\end{pmatrix}=:\bar{s},\quad\langle s,t\rangle=\bar{s}_{1}t_{1}+\bar{s}_{2}t_{2}.$ (120)
We extend complex conjugation to $\mathfrak{su}(2)$-valued spinors trivially, that is, for $s\in S$, $\upsilon\in\mathfrak{su}(2)$ we define $\overline{s\otimes\upsilon}=\overline{s}\otimes\upsilon$. Then under the isomorphisms (77), (79) the real structure on $(\Lambda^{1}\oplus\Lambda^{0})\otimes\mathbb{C}$ given by complex conjugation corresponds to $\nu\otimes\psi^{\ast}\mapsto\bar{\nu}\otimes\bar{\psi}^{\ast}$. Now the map $s\mapsto\bar{s}$ conjugates the Killing constant of any Killing spinor and conjugates the eigenvalue of any eigenspinor of $D^{A}-\operatorname{ad}(\Phi)$. So $s\mapsto\bar{s}$ gives antilinear maps $E_{x}^{\pm}\to E_{x}^{\mp}$ and $K^{\pm}\to K^{\mp}$. Therefore the two isomorphisms (113) are exchanged by the real structure on $T^{\mathbb{C}}_{x}M^{f}_{n,p}$.
Given a non-zero Killing spinor $\psi\in K^{-}$, we define an almost complex structure on $M^{f}_{n,p}$ by declaring that $T^{1,0}_{x}M^{f}_{n,p}=E_{x}^{+}\otimes\langle\bar{\psi}\rangle$ and $T^{0,1}_{x}M^{f}_{n,p}=E_{x}^{-}\otimes\langle{\psi}\rangle$, in which $\langle\psi\rangle\subset K^{-}$ denote the span of $\psi$. The associated linear map $J_{\langle\psi\rangle}:T_{x}M^{f}_{n,p}\to T_{x}M^{f}_{n,p}$ is then given by
$J_{\langle\psi\rangle}(\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi})=-i\nu\otimes\psi+i\bar{\nu}\otimes\bar{\psi}\quad\forall\nu\in E_{x}^{-}.$ (121)
To ensure that this is well-defined we must check that every real tangent vector can be written in the form $\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi}$, or in other words, that $\nu\mapsto\nu\otimes\psi+\bar{\nu}\otimes\bar{\psi}$ defines an isomorphism $E^{-}_{x}\to T_{x}M^{f}_{n,p}$. This map is injective because if $\nu$ is in its kernel then $0=\nu\otimes\psi^{\ast}(\bar{\psi})+\bar{\nu}\otimes\bar{\psi}^{\ast}(\bar{\psi})=-\langle\psi,\psi\rangle\nu$, and $\langle\psi,\psi\rangle\neq 0$ because $\psi\neq 0$. It is an isomorphism because its domain and target have the same real dimension.
These almost complex structures $J_{\langle\psi\rangle}$ coincide with the pluricomplex structures studied by Bielawski–Schwachhofer [5] and are therefore integrable. To see this, we write them out in local coordinates. We once again work in the upper half plane model and choose $\psi=\psi_{1}$ given in (109). Then a tangent vector $(a,\phi)$ can be written
$\operatorname{cl}_{S}(a,\phi)$ $=\begin{pmatrix}\phi-i\rho a_{\rho}&-i\rho(a_{x}-ia_{y})\\ -i\rho(a_{x}+ia_{y})&\phi+i\rho a_{\rho}\end{pmatrix}=\nu\otimes\psi^{\ast}+\bar{\nu}\otimes\bar{\psi}^{\ast},$ (122)
$\nu$ $:=-i\rho^{\frac{3}{2}}\begin{pmatrix}a_{x}-ia_{y}\\ -a_{\rho}+\frac{i}{\rho}\phi\end{pmatrix}.$
The almost complex structure (121) is therefore given by
$J_{\langle\psi\rangle}:(a_{x},a_{y},a_{\rho},\phi)\mapsto(-a_{y},a_{x},-\tfrac{1}{\rho}\phi,\rho a_{\rho}).$ (123)
In terms of the complex coordinates $z=x+iy$ and $v=\rho e^{i\theta}$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$, the components of $\nu$ are proportional to the holomorphic components $b_{z},b_{v}$ of the 1-form $b=a+\phi\,d\theta$. So $J_{\langle\psi\rangle}$ is the natural almost complex structure on the moduli space of circle-invariant instantons on $\mathbb{C}^{2}$, which is the same as the complex structure studied in [5].
It is worth pointing out that, although $J_{\langle\psi\rangle}$ was defined using the gauge fixing condition (25), the real tangent vector $a+\phi=\nu\otimes\psi^{\ast}+\bar{\nu}\otimes\bar{\psi}^{\ast}$ in (121) does not satisfy (25). This is because the two summands separately satisfy the two different gauge-fixing conditions in (25), and their sum satisfies neither. The most natural way to write the gauge-fixing condition satisfied by the real tangent vector in (121) is to regard the monopole as a circle-invariant instanton on $\mathbb{R}^{4}$. In terms of the 1-form $b=a+\phi\,d\theta$ on $\mathbb{R}^{4}$, the components $b^{1,0},b^{0,1}$ satisfy the two alternative gauge-fixing conditions (25). A calculation similar to (58) in section 3 shows that these can be written in terms of the Hodge star operator $*_{E}$ of the Euclidean metric on $\mathbb{R}^{4}$ as follows:
$d^{B}*_{E}\frac{1}{v^{2}}b^{1,0}=0=d^{B}*_{E}\frac{1}{\bar{v}^{2}}b^{0,1},$ (124)
where $v=\rho\mathrm{e}^{i\theta}$. Since $*_{E}b^{1,0}$ is a form of type $(2,1)$,
$d^{B}*_{E}\frac{1}{v^{2}}b^{1,0}=\bar{\partial}^{B}*_{E}\frac{1}{v^{2}}b^{1,0}=\frac{1}{v^{2}}\bar{\partial}^{B}*_{E}b^{1,0}=\frac{1}{v^{2}}d^{B}*_{E}b^{1,0}.$ (125)
This, together with a similar calculation for $b^{0,1}$, shows that the gauge-fixing condition (25) for $a+\phi$ is equivalent to $d^{B}*_{E}b=0$, which is the usual Coulomb gauge-fixing condition on $\mathbb{R}^{4}$. Written in terms of $(a,\phi)$ it reads $\partial^{A}_{x}a_{x}+\partial^{A}_{y}a_{y}+\partial^{A}_{\rho}a_{\rho}+a_{\rho}/\rho+[\Phi,\phi]/\rho^{2}=0$. This is clearly different from the gauge-fixing condition (25), which takes the form $\partial^{A}_{x}a_{x}+\partial^{A}_{y}a_{y}+\partial^{A}_{\rho}a_{\rho}+a_{\rho}/\rho+[\Phi,\phi]/\rho^{2}\pm 2i\phi=0$.
The complex structures $J_{\langle\psi\rangle}$ depend on a line $\langle\psi\rangle\in K^{-}$, or equivalently, a point in $\mathbb{P}K^{-}$. The space $\mathbb{P}K^{-}$ can be identified with the boundary $S^{2}_{\infty}$ in a natural way. For any non-zero Killing vector $\psi$, there is a unique point $z_{0}\in S^{2}_{\infty}$ such that $|\psi|$ decays as one approaches $z_{0}$ along any geodesic. Indeed, from the explicit solutions (109) we see that
$\bar{z}_{0}\psi_{1}+\psi_{2}=\rho^{-\frac{1}{2}}\begin{pmatrix}\bar{z}_{0}-\bar{z}\\ \rho\end{pmatrix}$ (126)
decays as $\rho^{\frac{1}{2}}$ as along geodesics approaching the point $z_{0}\in\mathbb{C}\subset S^{2}_{\infty}$, and grows as $\rho^{-\frac{1}{2}}$ along other geodesics. So there is a natural anti-holomorphic identification $S^{2}_{\infty}\cong\mathbb{P}K^{-}$, and a holomorphic identification $S^{2}_{\infty}\cong\mathbb{P}K^{+}$. Thus $\langle\psi\rangle\to J_{\langle\psi\rangle}$ is a map from $S^{2}_{\infty}$ to the space of complex structures on $T_{x}M^{f}_{n,p}$. This map is holomorphic, in the sense that $T^{1,0}_{x}M^{f}_{n,p}$ depends holomorphically on $z_{0}\in S^{2}_{\infty}$. This perspective on pluricomplex structures is the starting point of the paper [5].
Now let us consider the compatibility of the complex structures $J_{\langle\psi\rangle}$ with the bilinear forms $g_{\pm}$. Under the isomorphism (77), the complex-linear inner product on $(\Lambda^{1}\oplus\Lambda^{0})\otimes\mathbb{C}$ becomes $g(s\otimes t,s^{\prime}\otimes t^{\prime})=\frac{1}{2}(s,s^{\prime})(t,t^{\prime})$. So for any $\nu\otimes\psi\in E^{-}_{x}\otimes K^{-}$,
$g_{-}(\nu\otimes\psi,\nu\otimes\psi)=-\frac{1}{4}\int_{H^{3}}(\psi,\psi)\operatorname{Tr}\big{(}\nu,\nu\big{)}\mathrm{Vol}_{H^{3}}=0.$ (127)
So $g_{-}$ is compatible with the complex structures in the sense that it makes antiholomorphic tangent vectors null. Similarly, $g_{+}$ makes holomorphic tangent vectors null. If $g_{+}$ and $g_{-}$ were real this would imply that they are compatible with $J_{\langle\psi\rangle}$ in the usual sense that $g_{\pm}(J_{\langle\psi\rangle}X,J_{\langle\psi\rangle}X)=g_{\pm}(X,X)$ for all $X$.
We have seen that the tensor product decomposition (113) induces complex structures on the moduli space. A second interesting structure is a natural decomposition of the bilinear forms $g_{\pm}$. The Killing spinor equation implies that $(\psi_{1},\psi_{2})$ is a constant function on $H^{3}$ for any $\psi_{1},\psi_{2}\in K^{\pm}$. Therefore $g_{\pm}$ can be written
$g_{\pm}(\nu_{1}\otimes\psi_{1},\nu_{2}\otimes\psi_{2})=\frac{1}{2}(\psi_{1},\psi_{2})\omega_{\pm}(\nu_{1},\nu_{2}),\quad\forall\psi_{\alpha}\in K^{\pm},\nu_{\alpha}\in E_{x}^{\pm},$ (128)
in which
$\omega_{\pm}(\nu_{1},\nu_{2}):=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}\big{(}\nu_{1},\nu_{2}\big{)}\mathrm{Vol}_{H^{3}}.$ (129)
In other words, the $L^{2}$ symmetric form $g_{\pm}$ on $T_{x}M^{f}_{n,p}$ can be decomposed into an $L^{2}$ skew-symmetric form $\omega_{\pm}$ on $E_{x}^{\pm}$ and the pointwise symplectic pairing (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $K^{\pm}$. If $g_{\pm}$ is non-degenerate, then $\omega_{\pm}$ is also nondegenerate and is a symplectic form. So if $g_{\pm}$ is non-degenerate, the structure group $SO(4n)$ of the tangent bundle can be reduced to $Sp(2n)$. This $Sp(2n)$-structure is not expected to be integrable, and it would be interesting to determine its torsion classes.
From $\omega_{\pm}$, one can construct a family of holomorphic (symplectic) forms on the tangent bundle. These are parametrised by $\psi\in K^{\pm}$ and defined by
$\Omega_{\psi}^{\pm}((a_{1},\phi_{1}),(a_{2},\phi_{2})):=\omega_{\pm}((a_{1}+\phi_{1})\cdot\psi,(a_{2}+\phi_{2})\cdot\psi).$ (130)
With respect to the complex structures $J_{\langle\psi\rangle}$, $\Omega_{\psi}^{-}$ is of type $(2,0)$ and $\Omega_{\psi}^{+}$ is of type $(0,2)$.
The two-forms $\omega_{\pm}$ are reminiscent of a construction due to Nash [24]. To explain this, we first recall relevant part of [24]. Nash considered the minitwistor space, which is the space of oriented geodesics in $H^{3}$ and can be be identified with $S^{2}_{\infty}\times S^{2}_{\infty}$. From our perspective it is natural to identify the minitwistor space with $\mathbb{P}K^{+}\times\mathbb{P}K^{-}$. A monopole $x\in M^{f}_{n,p}$ is identified with a spectral curve $S$ in a bundle over minitwistor space, and the tangent space $T_{x}^{\mathbb{C}}M^{f}_{n,p}$ is identified with the space of sections of the normal bundle $\hat{N}$ of this curve. Nash obtains in eq. (28) of [24] two decompositions
$T_{x}^{\mathbb{C}}M^{f}_{n,p}$ $\cong H^{0}(S,\hat{N}(-1,0))\otimes H^{0}(S,\mathcal{O}(1,0))$ (131)
$T_{x}^{\mathbb{C}}M^{f}_{n,p}$ $\cong H^{0}(S,\hat{N}(0,-1))\otimes H^{0}(S,\mathcal{O}(0,1)).$ (132)
Here $\mathcal{O}(1,0)$ and $\mathcal{O}(0,1)$ are the duals of the tautological bundles of $\mathbb{P}K^{\pm}$; the spaces $H^{0}(S,\mathcal{O}(1,0))$, $H^{0}(S,\mathcal{O}(0,1))$ of sections are naturally identified with $K^{\pm}$. So Nash’s decompositions are very similar to (113). In Lemma 3.5 Nash constructs skew-symmetric pairings on his spaces $H^{0}(S,\hat{N}(-1,0))$, $H^{0}(S,\hat{N}(0,-1))$ and shows that they are non-degenerate. We conjecture that his decomposition coincides with (113), and that these symplectic forms coincide with (129). If true, this would imply that $\omega_{\pm}$ are nondegenerate and hence, via (128), that $g_{\pm}$ are nondegenerate.
5 Calculating the metric
In this section we obtain some explicit solutions of the gauge-fixing condition (25) and hence evaluate some components of the bilinear form $g_{\pm}$. In particular, we calculate $g_{\pm}$ for the 1-monopole moduli space and confirm that it is a metric.
Let $(A,\Phi)$ be a monopole of any charge and any mass parameter. Then
$(a,\phi)_{0}:=(d^{A}\Phi,0)$ (133)
solves the linearised Bogomolny equation (20) and both gauge-fixing conditions (25) because $(A,\Phi)$ solves the Bogomolny equation (3) and hence $d^{A}\ast d^{A}\Phi=d^{A}F^{A}=0$. The solution (133) is induced by the unframed gauge transformation $\Phi$. However, $\Phi$ is not a framed gauge transformation, because $\int_{H^{3}}\operatorname{Tr}(F^{A}\wedge d^{A}\Phi)=-4\pi np\neq 0$. So (133) is a non-zero tangent vector to $M^{f}_{n,p}$. Inserting (133) into (27) gives
$g_{\pm}((a,\phi)_{0},(a,\phi)_{0})=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(d^{A}\Phi\wedge\ast d^{A}\Phi)=E=2\pi np.$ (134)
From (133) and two independent Killing spinors $\psi_{1},\psi_{2}\in K^{\pm}$ we can construct two solutions
$\nu_{\alpha}=-d^{A}\Phi\cdot\psi_{\alpha}=F^{A}\cdot\psi_{\alpha}$ (135)
of the Dirac eigenspinor equation (87). Proposition 11 shows that these solve (87), and moreover that $\nu_{\alpha}\otimes\psi_{\beta}^{\ast}$ give four solutions of the linearised Bogomolny equation (20) and gauge-fixing condition (25). If $\psi_{1},\psi_{2}$ are chosen such that $(\psi_{1},\psi_{2})=1$ as in (109) then $\psi_{2}\otimes\psi_{1}^{\ast}-\psi_{1}\otimes\psi_{2}^{\ast}$ is the identity map on $S$ and so
$\nu_{1}\otimes\psi_{2}^{\ast}-\nu_{2}\otimes\psi_{1}^{\ast}=d^{A}\Phi$ (136)
corresponds to the tangent vector (133) that we started with. In this equation, and throughout this section, we implicitly make use of the isomorphism (77) between $\Lambda^{1}\oplus\Lambda^{0}$ and $S\otimes S^{\ast}$, and omit the notation $\operatorname{cl}_{H}$.
The remaining three tangent vectors can be interpreted using isometries. Given any Killing vector field $Z$, $(a,\phi)_{Z}:=(Z\lrcorner F^{A},Z\lrcorner d^{A}\Phi)$ differs from the Lie derivative $(L_{Z}A,L_{Z}\Phi)$ of the monopole $(A,\Phi)$ by a gauge transformation. So it describes a tangent vector to the moduli space induced by $Z$ and solves the linearised Bogomolny equation (20), but does not necessarily solve the gauge-fixing condition (25).
The isometry group of hyperbolic space is isomorphic to the Lorentz group $SO(3,1)$. This is generated by Killing vector fields $X_{i},Y_{j}$ satisfying
$[X_{i},X_{j}]=-\epsilon_{ijk}X_{k},\quad[X_{i},Y_{j}]=-\epsilon_{ijk}Y_{k},\quad[Y_{i},Y_{j}]=\epsilon_{ijk}X_{k}.$ (137)
In upper half-space coordinates they are given by
$X_{1}$ $=yY_{3}+\frac{1}{2}(1-|\mathbf{x}|^{2})\partial_{y},$ $Y_{1}$ $=-xY_{3}+\frac{1}{2}(1+|\mathbf{x}|^{2})\partial_{x},$ (138)
$X_{2}$ $=-xY_{3}-\frac{1}{2}(1-|\mathbf{x}|^{2})\partial_{x},$ $Y_{2}$ $=-yY_{3}+\frac{1}{2}(1+|\mathbf{x}|^{2})\partial_{y},$
$X_{3}$ $=x\partial_{y}-y\partial_{x},$ $Y_{3}$ $=x\partial_{x}+y\partial_{y}+\rho\partial_{\rho},$
where $|\mathbf{x}|^{2}=x^{2}+y^{2}+\rho^{2}$. We compare these with the vector fields $Z_{\psi_{\alpha},\psi_{\beta}}$ given by
$Z_{\psi_{\alpha},\psi_{\beta}}=-\psi_{\alpha}\otimes\psi_{\beta}^{\ast}-\psi_{\beta}\otimes\psi_{\alpha}^{\ast}$ (139)
under the isomorphism (77). For concreteness, we now choose the lower sign in the gauge-fixing condition (25) and choose $\psi_{1},\psi_{2}\in K^{-}$ as in (109). Using (105) one calculates e.g.
$Z_{\psi_{1},\psi_{2}}=\begin{pmatrix}1&2(x-iy)\rho^{-1}\\ 0&-1\end{pmatrix}=-X_{3}+iY_{3}.$ (140)
Proceeding similarly we find
$\begin{split}Z_{1}:=Y_{1}+iX_{1}&=i(\psi_{2}\otimes\psi_{2}^{\ast}-\psi_{1}\otimes\psi_{1}^{\ast}),\\ Z_{2}:=Y_{2}+iX_{2}&=-(\psi_{1}\otimes\psi_{1}^{\ast}+\psi_{2}\otimes\psi_{2}^{\ast}),\\ Z_{3}:=Y_{3}+iX_{3}&=i(\psi_{1}\otimes\psi_{2}^{\ast}+\psi_{2}\otimes\psi_{1}^{\ast}).\end{split}$ (141)
In general, the Clifford multiplication of the two-form $F^{A}$ with a vector field $Z$ is given by
$F^{A}\cdot Z=F^{A}\wedge Z-F^{A}\llcorner Z=(Z\lrcorner\ast F^{A})\mathrm{Vol}+Z\lrcorner F^{A}=Z\lrcorner d^{A}\Phi+Z\lrcorner F^{A},$ (142)
in which we used that $\mathrm{Vol}$ acts as identity in the Clifford algebra. Thus
$\begin{split}i(\nu_{2}\otimes\psi_{2}^{\ast}-\nu_{1}\otimes\psi_{1}^{\ast})&=(Z_{1}\lrcorner F^{A},Z_{1}\lrcorner d^{A}\Phi)=:(a,\phi)_{1},\\ -(\nu_{1}\otimes\psi_{1}^{\ast}+\nu_{2}\otimes\psi_{2}^{\ast})&=(Z_{2}\lrcorner F^{A},Z_{2}\lrcorner d^{A}\Phi)=:(a,\phi)_{2},\\ i(\nu_{1}\otimes\psi_{2}^{\ast}+\nu_{2}\otimes\psi_{1}^{\ast})&=(Z_{3}\lrcorner F^{A},Z_{3}\lrcorner d^{A}\Phi)=:(a,\phi)_{3}\end{split}$ (143)
are solutions of the linearised Bogomolny equation (20) and the gauge fixing condition (25). They are complexified tangent vectors and can be interpreted as complex linear combinations of boosts and rotations of the monopole.
We have now identified four complexified tangent vectors (136), (143), written in terms of $\nu_{\alpha}\in E_{x}^{-}$ and $\psi_{\alpha}\in K^{-}$. We can evaluate the bilinear form $g_{-}$ on them with the help of (128), which relates these to skew-symmetric pairings on $K^{-}$, $E_{x}^{-}$. Working in upper half space coordinates with the Clifford algebra representation (105) we get
$F^{A}=-i\rho^{2}(F_{xy}\sigma_{3}+F_{y\rho}\sigma_{1}+F_{\rho x}\sigma_{2})=-i\rho^{2}\begin{pmatrix}F_{xy}&F_{y\rho}-iF_{\rho x}\\ F_{y\rho}+iF_{\rho x}&-F_{xy}\end{pmatrix}.$ (144)
Then from (109) and (135) we compute
$\begin{split}\nu_{1}&=-i\rho^{3/2}\begin{pmatrix}F_{xy}\\ F_{y\rho}+iF_{\rho x}\end{pmatrix},\\ \nu_{2}&=i\rho^{3/2}\begin{pmatrix}(x-iy)F_{xy}-\rho(F_{y\rho}-iF_{\rho x})\\ (x-iy)(F_{y\rho}+iF_{\rho x})+\rho F_{xy}\end{pmatrix}.\end{split}$ (145)
Thus
$\begin{split}\omega_{-}(\nu_{1},\nu_{2})&=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(\nu_{1},\nu_{2})\operatorname{Vol}_{H^{3}}\\ &=-\frac{1}{2}\int_{H^{3}}\rho^{4}\operatorname{Tr}(F_{xy}^{2}+F_{y\rho}^{2}+F_{\rho x}^{2})\operatorname{Vol}_{H^{3}}\\ &=-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(F^{A}\wedge*F^{A})=2\pi np,\end{split}$ (146)
using (4). Thus, from (128),
$g_{-}(\nu_{\alpha}\otimes\psi_{\beta}^{\ast},\nu_{\gamma}\otimes\psi_{\delta}^{\ast})=\pi np\,\epsilon_{\alpha\gamma}\epsilon_{\beta\delta}$ (147)
It follows that
$g_{-}((a,\phi)_{\mu},(a,\phi)_{\nu})=2\pi np\,\delta_{\mu\nu},\quad\text{for}\,\mu,\nu\in\{0,1,2,3\}.$ (148)
Equation (148) is a satisfyingly simple formula for the symmetry-induced components of $g_{-}$. It is also consistent with an interpretation of charge $n$ monopoles as ensembles of $n$ point particles, each with mass $2\pi p$. To see this, we note that the Killing vector fields (138) satisfy
$g(Z_{j},Z_{k})=g(Y_{j},Y_{k})-g(X_{j},X_{k})+ig(X_{j},Y_{k})+ig(Y_{j},X_{k})=\delta_{jk}$ (149)
with respect to the metric $g$ on $H^{3}$. The motion of a particle of mass $m$ in $H^{3}$ is described by the metric $mg$. The motion of $n$ particles of equal mass is described by the product metric $g_{n}=mg+\ldots+mg$ on $(H^{3})^{n}$. So we find that
$g_{n}(Z_{j},Z_{k})=g_{n}(Y_{j}+iX_{j},Y_{k}+iX_{k})=nm\delta_{jk},$ (150)
in which $X_{j},Y_{j}$ represent simultaneous boosts and rotations of all $n$ particles. This agrees with our result (148) for the choice $m=2\pi p$. The total mass $2\pi np$ is equal to the energy $E$ of the monopole. It would be interesting to also compare inner products of real Killing vectors $X_{i},Y_{j}$ for the point particle and monopole metrics. However, we are not able to do this because we are currently only able to solve the gauge fixing condition (25) for the complex combinations $Y_{j}+iX_{j}$. For point particles, the inner products $g(X_{j},X_{k})$ etc depend on the position of the particle in $H^{3}$, so one would expect that for monopoles $g_{-}(X_{j},X_{k})$ depends on the moduli space coordinates, unlike the inner products (148).
Equation (148) fully determines $g_{-}$ in the case of charge 1. It is well-known that the Bogmolny equations admit a spherically-symmetric solution when $n=1$. The moduli space $M^{f}_{1,p}$ can be generated by transforming this monopole using isometries of $H^{3}$ and unframed gauge transformations $g=\exp(s\Phi)$. It is therefore diffeomorphic to ×$\mathbb{R}\times SO(3,1)/SO(3)=\mathbb{R}\times H^{3}$. The bilinear form $g_{-}$ is clearly invariant under the transitive action of $SO(3,1)\times\mathbb{R}$, so it is determined by its value at a single point in the moduli space.
Consider the point in the moduli space corresponding to the spherically-symmetric 1-monopole. This is identified with the point $(x,y,\rho)=(0,0,1)$ in $H^{3}$. Rotation of this monopole is equivalent to a gauge transformation, so
$(X_{j}\lrcorner F^{A},X_{j}\lrcorner d^{A}\Phi)=(d^{A}\chi_{j},[\Phi,\chi_{j}])$ (151)
for some gauge transformations $\chi_{j}$. These gauge transformations are framed, because
$\int_{H^{3}}\operatorname{Tr}(d^{A}\chi_{j}\wedge\ast d^{A}\Phi)=\int_{H^{3}}\operatorname{Tr}(X_{j}\lrcorner F^{A}\wedge F^{A})\\ =\frac{1}{2}\int_{H^{3}}X_{j}\lrcorner\operatorname{Tr}(F^{A}\wedge F^{A})=\int_{H^{3}}0=0.$ (152)
So
$(a,\phi)_{j}=(Z_{j}\lrcorner F^{A},Z_{j}\lrcorner d^{A}\Phi)=(Y_{j}\lrcorner F^{A}+id^{A}\chi_{j},Y_{j}\lrcorner d^{A}\Phi+i[\Phi,\chi_{j}])$ (153)
and the tangent vectors $(a,\phi)_{j}$ represent boosts of the monopole. They can therefore be identified with the tangent vectors $Y_{j}$ given in (138), evaluated at the point $(0,0,1)\in H^{3}$. By direct calculation, $g_{H^{3}}(Y_{j},Y_{k})=\delta_{jk}$ at this point. Comparing with (148), we conclude that the metric on $M^{f}_{1,p}$ is
$g_{-}=2\pi p(g_{\mathbb{R}}+g_{H^{3}}).$ (154)
Since this is real it agrees with $g_{+}$.
In the case $p=\frac{1}{2}$ the result (154) for the metric on $M^{f}_{1,\frac{1}{2}}$ can also be verified by a direct calculation based on explicit formulae for the monopole. We include these calculations here as they provide a consistency check on our results, and they might be useful in future calculations for $n>1$. A monopole in $M^{f}_{1,\frac{1}{2}}$ corresponds to a circle-invariant instanton with $c_{2}=2np=1$. By writing the instanton using the ’t Hooft ansatz and changing to a circle-invariant gauge we obtain:
$\begin{split}A=\frac{i}{\lambda^{2}+|\mathbf{x}-\mathbf{x}_{0}|^{2}}\Big{[}&(\rho\sigma_{2}-(y-y_{0})\sigma_{3})dx+((x-x_{0})\sigma_{3}-\rho\sigma_{1})dy\\ &+((y-y_{0})\sigma_{1}-(x-x_{0})\sigma_{2})d\rho\Big{]},\end{split}$ (155)
$\Phi=\frac{i}{2}\sigma_{3}-\frac{i\rho}{\lambda^{2}+|\mathbf{x}-\mathbf{x}_{0}|^{2}}(\rho\sigma_{3}+(x-x_{0})\sigma_{1}+(y-y_{0})\sigma_{2}),$ (156)
in which $x_{0},y_{0},\lambda$ parametrise the monopole centre with respect to upper half space coordinates on $H^{3}$ and $|\mathbf{x}-\mathbf{x}_{0}|^{2}=(x-x_{0})^{2}+(y-y_{0})^{2}+\rho^{2}$. The corresponding field strength is
$F_{xy}=if\sigma_{3},\quad F_{y\rho}=if\sigma_{1},\quad F_{\rho x}=if\sigma_{2},$ (157)
with
$f=\frac{2\lambda^{2}}{(|\mathbf{x}-\mathbf{x}_{0}|^{2}+\lambda^{2})^{2}}.$ (158)
Using (145) one calculates
$\operatorname{Tr}(\nu_{1},\nu_{2})=-6\rho^{4}f^{2}.$ (159)
Therefore
$-\frac{1}{2}\int_{H^{3}}\operatorname{Tr}(\nu_{1},\nu_{2})\operatorname{Vol}_{H^{3}}=\\ \int_{H^{3}}\frac{12\lambda^{4}\rho^{4}}{((x-x_{0})^{2}+(y-y_{0})^{2}+\rho^{2}+\lambda^{2})^{4}}\frac{dx\,dy\,d\rho}{\rho^{3}}=\pi,$ (160)
which agrees with (146) since $n=1$ and $p=\tfrac{1}{2}$. The metric $g_{-}$ is therefore given by (147).
We can also obtain (153) by a direct calculation. To do so, we must choose $\lambda=1$, $x_{0}=y_{0}=0$ and find functions $\chi_{j}$ such that
$(a,\phi)=(Y_{j}\lrcorner F^{A},Y_{j}\lrcorner d^{A}\Phi)+i(d_{A}\chi_{j},[\Phi,\chi_{j}])$ (161)
is equal to a linear combination of $\nu_{\alpha}\otimes\psi_{\beta}$. For example, take $j=1$. One has
$Y_{1}\lrcorner F^{A}$ $=\frac{i}{(1+|\mathbf{x}|^{2})^{2}}\Big{[}(2x\rho\sigma_{1}+(1-x^{2}+y^{2}+\rho^{2})\sigma_{3})dy$
$-(2xy\sigma_{1}+(1-x^{2}+y^{2}+\rho^{2})\sigma_{2})d\rho+2x(y\sigma_{3}-\rho\sigma_{2})dx\Big{]},$ (162)
$Y_{1}\lrcorner d^{A}\Phi$ $=\frac{i\rho}{(1+|\mathbf{x}|^{2})^{2}}[2x(y\sigma_{2}+\rho\sigma_{3})-(1-x^{2}+y^{2}+\rho^{2})\sigma_{1}].$ (163)
It is now useful to note that
$\operatorname{Tr}\left[\begin{pmatrix}\Phi&0\\ 0&\Phi\end{pmatrix}(\phi+a)\right]=2\operatorname{Tr}(\Phi\phi)$ (164)
is a gauge invariant quantity. For $(a,\phi)=(Y_{1}\lrcorner F^{A},Y_{1}\lrcorner d^{A}\Phi)$ we find using (156) and (163) that
$2\operatorname{Tr}(\Phi\phi)=-\frac{8x\rho^{2}}{(1+|\mathbf{x}|^{2})^{3}}.$ (165)
We also have
$G_{1}:=2\operatorname{Tr}\left[\nu_{1}\otimes\psi_{1}^{\ast}\begin{pmatrix}\Phi&0\\ 0&\Phi\end{pmatrix}\right]$ $=-\frac{4i\rho^{2}(x+iy)}{(1+|\mathbf{x}|^{2})^{3}},$ (166)
$G_{2}:=2\operatorname{Tr}\left[\nu_{1}\otimes\psi_{2}^{\ast}\begin{pmatrix}\Phi&0\\ 0&\Phi\end{pmatrix}\right]$ $=2\operatorname{Tr}\left[\nu_{2}\otimes\psi_{1}^{\ast}\begin{pmatrix}\Phi&0\\ 0&\Phi\end{pmatrix}\right]=\frac{2i\rho^{2}(|\mathbf{x}|^{2}-1)}{(1+|\mathbf{x}|^{2})^{3}},$ (167)
$G_{3}:=2\operatorname{Tr}\left[\nu_{2}\otimes\psi_{2}^{\ast}\begin{pmatrix}\Phi&0\\ 0&\Phi\end{pmatrix}\right]$ $=\frac{4i\rho^{2}(x-iy)}{(1+|\mathbf{x}|^{2})^{3}}.$ (168)
Since
$i(G_{3}-G_{1})=2\operatorname{Tr}(\Phi\phi),$ (169)
it must be that (5), (163) corresponds to $i\nu_{2}\otimes\psi_{2}^{\ast}-i\nu_{1}\otimes\psi_{1}^{\ast}$, in agreement with (143). We therefore look for a function $\chi_{1}$ satisfying
$(Y_{1}\lrcorner F^{A},Y_{1}\lrcorner d^{A}\Phi)+i(d_{A}\chi_{1},[\Phi,\chi_{1}])=i(\nu_{2}\otimes\psi_{2}^{\ast}-\nu_{1}\otimes\psi_{1}^{\ast}).$ (170)
A computation shows that (170) is equivalent to the conditions
$i[\Phi,\chi_{1}]$ $=\frac{\rho(2\rho y\sigma_{3}+2xy\sigma_{1}+(1-x^{2}+y^{2}-\rho^{2})\sigma_{2})}{(1+|\mathbf{x}|^{2})^{2}},$ (171)
$id^{A}\chi_{1}$ $=\frac{1}{(1+|\mathbf{x}|^{2})^{2}}\big{[}d\rho\left(\sigma_{1}\left(\rho^{2}+x^{2}-y^{2}-1\right)+2xy\sigma_{2}\right)$
$+2ydy\left(\rho\sigma_{1}-x\sigma_{3}\right)-dx\left(\sigma_{3}\left(\rho^{2}+x^{2}-y^{2}-1\right)+2\rho y\sigma_{2}\right)\big{]}.$ (172)
The above equations have the unique solution
$\chi_{1}=i\frac{\rho\sigma_{1}-x\sigma_{3}}{1+|\mathbf{x}|^{2}}.$ (173)
With similar computations we find that
$(Y_{2}\lrcorner F^{A},Y_{2}\lrcorner d^{A}\Phi)+i(d_{A}\chi_{2},[\Phi,\chi_{2}])$ $=-(\nu_{1}\otimes\psi_{1}^{\ast}+\nu_{2}\otimes\psi_{2}^{\ast}),$ (174)
$(Y_{3}\lrcorner F^{A},Y_{3}\lrcorner d^{A}\Phi)+i(d_{A}\chi_{3},[\Phi,\chi_{3}])$ $=i(\nu_{1}\otimes\psi_{2}^{\ast}+\nu_{2}\otimes\psi_{1}^{\ast}),$ (175)
with
$\chi_{2}=i\frac{\rho\sigma_{2}-y\sigma_{3}}{1+|\mathbf{x}|^{2}},\quad\chi_{3}=\frac{i\sigma_{3}}{1+|\mathbf{x}|^{2}}+\Phi,$ (176)
which confirms the result (143), (153).
Appendix A Alternative definitions of the framed moduli space
Our definition of the framed moduli space $M^{f}_{n,p}$ differs from others that can be found in the literature. Atiyah–Hitchin [2] defined framed gauge transformations to be those that equal the identity at a chosen point on $S^{2}_{\infty}$. This is convenient for establishing a correspondence with rational maps, but unlike our definition it breaks some symmetry. Atiyah–Hitchin’s definition leads to a framed moduli space that is a circle-bundle over the unframed moduli space. The circle action is given by unframed gauge transformations $g_{s}=\exp(s\Phi)$, and these act trivially if and only if $s\in\frac{2\pi}{p}\mathbb{Z}$. This circle bundle admits a flat connection defined using the resultant of the associated rational map. The horizontal directions are those which leave the phase of the resultant of this map unchanged. This flat bundle is canonically identified with
$\tilde{M}^{u}_{n,p}\times_{\mathbb{Z}_{2}}S^{1}$ (177)
in which $\tilde{M}^{u}_{n,p}$ is a double-cover of $M^{u}_{n,p}$.
The $\mathbb{Z}_{2}$ action on the fibre $S^{1}$ of (177) is generated by the gauge transformation $g=-I_{2}$. This acts trivially on our framed moduli space $M^{f}_{n,p}$, so it is natural to compare our moduli space with the $\mathbb{Z}_{2}$-quotient of (177):
$(\tilde{M}^{u}_{n,p}\times_{\mathbb{Z}_{2}}S^{1})/\mathbb{Z}_{2}=\tilde{M}^{u}_{n,p}/\mathbb{Z}_{2}\times S^{1}/\mathbb{Z}_{2}\cong{M}^{u}_{n,p}\times S^{1}.$ (178)
This is trivial circle bundle equipped with a flat connection. The circle action is again generated by $g_{s}=\exp(s\Phi)$, but now these act trivially if and only if $s\in\frac{\pi}{p}\mathbb{Z}$.
The gauge transformations $\exp(\frac{\pi}{p}\Phi)$ act non-trivially on our framed moduli space $M^{f}_{n,p}$. So to compare with (178) we need to quotient by this action. From (15), the quotient is
$M^{f}_{n,p}/\mathbb{Z}=M^{u}_{n,p}\times S^{1},$ (179)
which is again a trivial circle bundle with a flat connection. The horizontal directions for the flat connection are those which do not change the Chern-Simons number. The flat connections give a canonical identification of (178) and (179).
To summarise the discussion: our framed moduli space is different from the framed moduli space of (177). But a $\mathbb{Z}$-quotient of ours is canonically identified with a $\mathbb{Z}_{2}$-quotient of the latter.